A Bibliography of the Use of TEX

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 May 2024 Version 2.17

Abstract

This bibliography records books and periodicals that have been published using TEX.

Title word cross-reference

-RÉ [Bau89].

1699-1777 [ML91]. 1929-1936 [FDK^{+}86]. 1980s [GFOE87, GP87].

2 [TNY94]. 20th [Pet92b, Stu16]. 25th [Ras91].

3 [She92].

4.x [VMS88].

5 [Lim10]. 5-th [Lim10]. 5.x [VMS88].

9 [Dib92]. ’90 [C02te].

A.I. [WP84]. Abaci [Sig02]. Abstract [AK91]. Access [C61te]. accounts [ML91]. Accuracy [Hig96]. Acne [C80te]. Acquisition [Ber85, BW84]. Action [Cha91]. ad [Mil96]. Ada [WWF87]. Adaptive [ÅW89, Pao89]. Advanced [BB90, BBBH92, Dib92, KN94, Sal95, Pic92]. Advances [BB89, BB90]. Africa [C03te]. Age [RR92]. Aged [C14te]. ai [RFB91]. Aided [LS89, Tay91]. AIDS [C01te, C02te, C04te, C03te]. ailed [Vol92]. Air [Prz91]. Aircraft [His96]. Alert [C10te, C09te, C11te, C15te, C34te, C16te, C61te, C70te, C79te]. Algebra [Cuo89, DW89a, DW89b, GT88a, GMW91, HWH91, Sha90, MW91]. Algebraic [MW91, HW91]. Algebras [BdK90]. Algérie [Ham90]. Algorithmic [MR91]. Algorithms [CLR89, Dev86, GK82, GK81, Hig96, Knu97, Knu98a, Knu00, Knu08, Lyn96, Omo94, PC89, PS93, Sed88, Sed83, Sed90, Wil86]. Almond [ACC^{+}95]. ämbetsman [And90]. Ambulatory [C05te]. America [Pet92b]. American [ABS88, Alf92, AMS8x]. Amerike [Pet92b]. amtematica [Pav93]. Anaesthesia [LL87]. Anaesthesiology [C40te]. Anaesthetist [C06te]. Analisi [Pav92, Boi90, CV90, Pav93]. analitica [Boi90]. analiz [Vol92]. Analog [Mea89]. Analysis [BB90, BBBH92, Dev90, GK82, GK81, Knu00, Mun91, New88, Pav92, Sch91, Var91, Vit87, Wei90a, Wei90b, Wel90, Pav93, Vol92]. Ancient [Tho88]. Angina [C25te]. anglijski [Sos92]. Anniversary [Gra88, Ras91]. Annual [C0790, C0890, C7389, C7490]. Anti [C11te]. Anti-inflammatory [C11te]. Anticancer [C10te]. antichnojj [GKL^{+}92]. Antimicrobial [C09te]. Antique [GKL^{+}92]. appendixes [ML91]. Application [JT98, L^{+}92]. Applications [BP89a, BB90, BBBH92, Fie89, GT88b, Taj89, BdK90]. Applied [Hoe90, Kee88, OD91, Wig90]. Apprentice [RW90]. Approach [DW89a, DW89b, DW89c, Kam90, MG89, Var92, Wer91, DW90b]. Arab [Con89]. Arbeitsbuch [LHR^{+}91a, LHR^{+}91b]. Archaeology [RR92]. Architecture [App91a, App90, App91b, Dal87, JA90, Omo94, ST91, Wak89]. Argument [Ash90]. Arithmetic [Omo94, WF82]. Art [Ham91, Keh93, Knu05a, Knu05b, Knu08, Knu09, PFTV86, SF89, SS86]. Artery [C35te]. Arthritis [C33te]. Articles [L^{+}92]. Artificial [Aya91, DL82, Lan89, WB79, Win89b, WP84, WS90a, WS90b]. artists [HH93]. Asian [BS91]. Aspects [LL87]. Assessment [C36te, Gib90]. Asthma [C17te, C26te]. Astrodynamics [BA96]. Astrophysics [Taj89]. Atheroma [C12te].

Atherosclerosis [C12te]. Atlante [TP90a]. Atlas [C13te, C14te, C8390, C8491, C8590, C8691, C8190, C8291, Ham90, TP90b, TP91]. Attack [Smi90]. attivi [Bie87]. attraverso [Boi90]. Audio [Ram98]. Authors [Fly95]. Automata [Gut90]. Automatic [AK93]. automatico [OP88]. avanzate [Pic92]. Axioms [Knu92].

Balance [Rai86]. Banners [Win89a]. base [Pic92]. Based [AAH88, Bar90, CRR^{+}90, DL82, Kam90, TWW88, Wer91]. Basic [Pie91, Kup87]. Basis [BW84, MW90a, MW90b]. Bawerk [And90]. Beauty [FvGGM90]. Beginner [Keh93]. Behavioral [MC88]. Beyond [GP87]. Bible [Knu91, She92]. Bibliography [ABS88, GFOE87]. binary [BvdH91, Knu09]. Biology [C37te, C43te, C60te, OALD97]. Biophysics [NL90]. Biotechnology [C15te, C41te]. birthday [FvGGM90]. Bitwise [Knu09]. Black [GKL^{+}92]. Blood [C05te, C64te, C68te]. Body [SJ87]. Böhm [And90]. Böhm-Bawerk [And90]. bolen [Vol92]. Book [Gra90a, Ros91, She92, Sig02]. bookbinders [HH93]. Booklet [PDG90]. booksellers [HH93]. Boolean [Knu08]. Borgerskapets [And90]. born [Stu16]. Both [Str92]. Bound [Dai92]. Bowyer [ML91]. Brazilian [MW91]. Bucket [Dev86]. Buheiry [Con89]. Building [TNY92, TNY94]. Business [WP84, FvGGM90]. byl [Abr91, Vol92].

C

[JCL89, FH95, Jae91, Jae89, KR99, Sed90]. Calculation [Sig02]. Calculus [Ada89]. California [ACC^{+}95]. Campi [SV91]. Cardiac [C0790]. Cardiology [C21te, C42te]. Cardiovascular [C16te, C38te]. Carolinae [CMU89]. Case [OALD97]. Castellano [W^{+}86]. Catalogue [Pal89, Win89a, WAD16]. Category [BW90, Pie91]. Cell [C43te, NL90, OALD97]. Cellular [Gut90]. Century [Lim10, Pet92a, Pet92b, Stu16]. Chaos [PJS92, PJS93]. Chaotic [Dev87, Dev89, PC89]. checklist [ML91]. Chem [Vol92]. Chinese [Lim10]. Chlorine [AAdR85]. Cientìfica [W^{+}86]. Circuit [Gir90]. Circuit-Switched [Gir90]. Circuiti [Bie87, Bec88, RFB91]. Circuits [Bak90, Sie86]. City [HH93]. Classes [MS88]. Classic [C17te, C18te, C19te, C20te]. Clausal [Ric89]. Clinical [C04te]. Clinician [C21te, C22te, C23te, C24te, C25te, C26te, C27te, C28te, C29te, C30te, C31te, C32te, C33te]. CMU [Ras91]. CNS [C34te]. Coalesced [Vit87]. Coast [GKL^{+}92]. Coastal [Cla84]. Coding [Sal05a]. Collected [FDK^{+}86]. Collection [Str92, WAD16]. College [DW89a, DW89b]. Combinations [Knu05b]. Combinatorial [Aig88, Knu08]. Combinatorics [vLW92]. Commemorative [Ras91]. commentary [ML91]. Commentationes [CMU89]. Commercial [WP84]. Common [Bro85, Ste90a]. Communication [Hoa91]. Communications [Dai91, O’S87, Sal05a]. Commutative [Sha90]. Comparative [Wel90]. Compilateurs [ASU89]. Compiler [FH95]. Compilers [ASU89]. complementi [Pav92]. complements [Pav92]. Complete [Sal98, Sal00, Sal04, Sal07]. Complex [AAP88, Jen90, CS87]. Complexity [BBR87, Par93, Jud90, TWW88, Wer91, Wil86, Zur90]. Compression [Sal98, Sal00, Sal04, Sal07, Sal01]. Computation [HPK91, LP92, Thi88, WH90, NC89]. Computational [BBR87, All91, Gri81, KM90, Taj89, Van92, Wer91]. Computations [GV89]. Computer [AwJS85, Ang89, BW90, Dai91, GT88a, GT88b, Gri91, HS91, HWH91, Har85, Har86, Har87, Kel82, Knu05a, Knu05b, Knu08, Knu09, KS91, KN94, LS89, MW91, MW90a, MW90b, Omo94, Pie91, Ras91, Rog85, Sal99, Sal06a, Sal06b, SB90, SBK89, Tay91, Wol88, Wol90a, JCo92, Bie91, Pic92, Sal05a, Sal05b]. Computer-Aided [Tay91]. Computers [BM90, DDSv91, SHH^{+}89, Pic92]. Computing [AK93, Big89, FWM94, GH93, GO92, MW91, Nas90, PFTV86, Thi88, Wat90]. con [Boi90]. concept [Bie91]. Concepts [BP89b]. Concise [Wei99]. Concordantia [Mil96]. Concrete [GKP89, GKP94, MG89]. Concurrency [Hoa91]. Concurrent [Dal87]. Condensed [RS88, Tes90]. Connection [Hil85]. Constraints [Gri91]. Contemporary [Dai92]. Content [Ber93]. Control [AAH88, ÅW89, BHJ^{+}82, Cra89, FPW90, FPEN91, Pau81]. Cookbook [Stu01]. Cornishman [Corte]. Coronary [C13te, C35te, C36te]. corrected [PJS93]. Correlated [Fra91]. Cosmic [Sok89]. Counselling [C04te]. Course [Kel82, Kre90, SB90, vLW92]. Craft [O’K90]. Cray [RR89]. CRC [Wei99]. Critical [Prz91]. Culture [ES90, GKL^{+}92, L^{+}92]. Cumulative [Pal89]. Current [C37te, C38te, C39te, C40te, C41te, C42te, C43te, C44te, C45te, C46te, C47te, C48te, C49te, C50te, C51te, C52te, C53te, C54te, C55te, C56te, C57te, C58te, C59te, C60te, C61te, C62te, C63te, C8390, C8491, C8590, C8691]. Curves [Sal06a].

Dam [DE90]. Data

[Dal87, KR99, Sal98, Sal00, Sal03, Sal04, Sal07, Wei90b, Sal01, Sal05a]. Database [BB89, Lus88, Wie87]. Databases [PSF91]. December [Wel89b]. Decision [Knu09]. Defense [Prz91]. del [Pic92, TP90a]. Dentistry [C44te]. d’esame [SV91]. Descriptive [WAD16]. Design [CM88, CRR^{+}90, FH95, GC94, His96, LS89, Moy95, PS93, Jud90, Tay91, Vit87, Wie87, KR99]. Designers [WF82]. Desk [wJBH86]. Detection [C64te, Sch91]. Development [C46te, Dij89, JT98, Jon90]. Developments [Hoa91]. Diabetes [C71te]. Diagrams [Knu09]. Dialogue [Car90]. Dictionary [C21te, C22te, C23te, C24te, Jae91, Rw91, Nic93]. Differential [Bir91, DFH^{+}90, GO92, Hel90, Wer91, Zwi89, HW91]. differential-algebraic [HW91]. Differentsial’nye [Bir91]. Digital [FPW90, Uli87, WF82]. digitali [RFB91]. Dijkstra [FvGGM90]. dimension [BdK90]. Dimensioning [Gir90]. Discovery [PSF91]. Discrete [CDF89, MR91, Pif91]. Disease [C12te, C13te, C14te, C35te]. Diseases [C48te]. disegno [OP88]. Disordered [HR90]. Disorders [C64te]. Distance [BH90]. Distributed [Lyn96]. DNA [BM90]. Document [App91a, App90, App91b]. documentary [Rau92]. Doing [Wol88, Wol90a]. Dokumentaustausch [App90]. Dokumentenverbeitung [BK89]. Drawings [Win89a]. Drugs [C38te, C39te]. Dynamic [FPW90, FPEN91, MC88]. Dynamical [Dev87, Dev89, Wig90].

Early [Gri81, KT90, Stu16]. East [ABS88]. Eco [RBM^{+}91]. Eco-Logic [RBM^{+}91]. Ecology [Lea85, MC88, OALD97]. Econometrics [ACC^{+}95]. Economic [Var93]. Economies [Cas90]. Economy [AAP88]. Edition [KN94]. Edsger [FvGGM90]. Ehtudy [GKL^{+}92]. Eichtheorien [Sch88]. Einführung [App90, BK89]. Elderly [C66te, C72te]. Electrical [IEE89, SB90]. Electrodynamics [Par91, Dou90]. Electromagnetics [Sad92]. Electron [Fra91]. Electronic [Sol92]. Electronics [KSV91]. Elektrotehniski [EVxx]. Elementary [Kan88]. Elements [Kup87, LP92, Rog85, SJ87, Thi88]. elettriche [BB88]. elettromagnetici [SV91]. Elites [MS88]. Encapsulation [Gri89]. Encyclopedia [Wei99]. Endoscopy [C0890, C8190, C8291]. Energy [Sok89]. Engineering [DE90, KN94, Poz90, Rob95, SB90, WL89]. Engineering. [Kah98]. Engineers [Ham91, Pif91]. English [Bir91, GKL^{+}92, Pav92, Pet92a, Ste92, Sos92, ASU89, Pav93, Sig02]. English-language [Sos92]. engravers [HH93]. Ensam [Ber89]. Entropy [Zur90]. Entwicklung [Kah98]. Eólico [TP90b]. Éolien [TP91]. Equations [Bir91, DFH^{+}90, GO92, Wer91, Zwi89, HW91]. Erwin [KN94]. esame [Pav93]. Esercitazioni [BB88]. Esercizi [BV90a, Pav92]. Español [Bau89]. Essays [ES90, MS88]. Ester [Ber89]. Estimation [Sch91]. Eugen [And90]. Europäischer [TP90c]. European [ABS88, C65te]. Européen [TP91]. Europeo [TP90b, TP90a]. Everyday [Stu01]. Evolution [PK91, BvdH91]. Evolving [AAP88]. examination [Pav93]. Exercises [Pav92]. Exhibition [Wel89b, Wel89a]. Experience [BP89a]. Experiment [Gut90]. Experimentalphysiker [Sch88]. Experiments [MB90]. Expert [Lus88]. Explained [GC94]. Export [ACC^{+}95].

fågel [Ber89]. Failure [C27te, C67te].

family [Wak89]. Fascicle

[Knu05a, Knu05b, Knu08, Knu09]. Fast [C10te, C09te, C11te, C15te, C34te, C16te, C61te, C70te, C79te, Van92]. Fast-Alert [C10te, C09te, C11te, C15te, C34te, C16te, C61te, C70te, C79te]. father [ML91]. Feedback [FPEN91]. Feet [Dai92]. Fertilizer [C79te]. Festkörpers [HwCZ91]. Feynman [Sch88]. Feynman-Graphen [Sch88]. Fibonacci [Sig02]. Field [Hat91, Ram90, Tes90]. File [Wie87]. Financial [Var93]. Finite [BdK90, Hoe90].

Firewalls [CBR03]. First [CGG^{+}92, Kel82, Pal89, SB90, MW91]. Firstbook [Ste92]. fizik [Abr91]. Fontographer [Moy95]. förbindelser [Hol88]. Force [Prz91]. Forest [Lea85]. Form [Ric89]. Formal [Dij89]. Format [App90, App91b]. Formation [BvdH91, Con89]. Fortieth [Gra88]. Fortran [Cla86, JC88]. Foundation [CM88]. Foundations [Hue89, LMW88, OD91, Sal05b, vN18]. Fourier [Van92, BB90]. Fractals [PJS92, PJS93]. Fracture [HR90]. Frameworks [Van92]. Freezing [HLZJ91]. Friday [Wel89b, Wel89a]. Friend [C06te]. Frontiers [PJS92, PJS93]. Functional [Dev90, Hue89, Tho91, Tur90]. Functioning [Syz91]. Functions [DW90a, Knu08, Pav92]. Fundamental [Knu97]. Fundamentals [BA96, Cal88, CDF89]. Fungicide [C70te]. funzionale [CV90]. funzioni [Pav92]. Fusion [Taj89]. Future [SBK89]. Futuro [Bau89].

Garden [GC94]. Gas [DFH^{+}90]. Gastroenterology [C22te, C45te, C65te, C66te]. Gastrointestinal [C0890, C8190, C8291]. gde [Abr91]. General [HvM92]. Generating [Knu05b]. Genetic [Wei90b]. Genetics [C46te]. Geography [Sol92]. geometria [Boi90]. Geometric [Gri91, LS89, Sal99]. Geometry [Ad81, Hel90]. Geophysical [DTZJ90]. Get [Gre91]. Glass [HLZJ91]. Gödel [Hil94, FDK^{+}86]. Gogol [Rau92]. Government [Smi90]. Graduale [Mil96]. Grammatical [BW84]. Graphen [Sch88]. Graphics [Ang89, Rog85, Sal99, Sal06a, Sal06b]. Graphing [DW89a, DW89b, DW89c, DW90b]. Graphs [BH90, DW90a]. gravitation [Dou90]. Greece [Tho88]. Greek [WAD16].

Groundwater [Str89]. Growth [C70te]. Guide [ABB^{+}92, C04te, C80te, Dib92, Fly95, GHKD90, HWH91, Keh93, Lus88, Mat85, TNY92, TNY94, Sal01, Seh92, TVG87]. Gynecology [C52te].

Hacker [CBR03, Rw91]. Halftoning [Uli87]. Handbook [C67te, C68te, C69te, Fly95, For89, Kra01, Zwi92, Zwi89]. handedly [Ros90]. Hands [MS85]. Harbor [Smi90]. Hartley [Bra86]. Hashing [PS93, Vit87]. Health [Gra88]. Hearings [Smi90]. Heart [C12te, C13te, C14te, C27te, C67te]. Heavy [BW91]. Heisenberg [dB90]. Hemindustrin [Ahl88]. Henning [Ber89]. Hepatology [C65te]. Herbicide [C70te]. Higgs [GHKD90]. High [BCH^{+}90, C68te, Fra91, Hal88]. Hindu [Les91]. Historical [BS91, LL87, Lim10, Nic93]. History [GKL^{+}92, MS88, Nas90, Pal89, Stu16, Win89a, wJBH86]. Honoring [Van89]. Hormone [C10te]. Horsmonden [Stu16]. hospital [Ber89]. HTML [Fly95]. Hulls [Knu92]. Human [Gri81]. Humanities [Rah87]. Hunter [GHKD90]. Hyperlipidaemia [C18te, C28te, C69te]. Hypertension [C19te, C23te, C29te, C7389, C7490, C71te, C72te, C76te, C77te].

Idrott [Hol88]. idrottsrörelsens [Hol88]. IEEE [IEE89]. II [HW91, HLZJ91, Wel89a, Wol89]. Illuminated [Knu91, She92]. Illustrated [C21te, C22te, C23te, C24te]. Image [Ber93]. Images [Gri81]. Immune [PK91]. Immunobiology [C75te]. Immunology [C47te, Per88a, Per88b]. Impatient [AL91]. Implementation [FH95, Omo94, Syz91]. Improper [Pav92]. impropri [Pav92]. Index [Smi90]. indexes [ML91]. Industry [ACC^{+}95]. Infarction [C25te]. Infectious [C48te]. infinite [BdK90]. Infinitivo [Bau89]. inflammatory [C11te]. informacji [Bie91]. Information [App91b, Rah87, RR92, Wer91, Zur90, TWW88, Bie91]. Information-Based [Wer91, TWW88]. Insecticide [C79te]. Insights [Dib92]. Institute [Prz91, Pal89, Win89a, wJBH86, BB90, BBBH92]. Instruction [Cha91]. instrumenta [Mil96]. Instrumentation [KS91, SBK89]. Insulation [IEE89]. Integral [Wer91, Pav92]. Integrali [Pav92]. integrals [Pav92]. integrati [RFB91]. Integration [CW83, Zwi92]. Intelligence [DL82, WB79, Win89b, WP84, WS90a, WS90b]. Intelligent [ST91]. Interaction [Lea85, SJ87, Dou90]. Interactions [Kru88]. Interchange [App90, App91b]. Interconnection [JA90]. Interconnections [Bak90]. Interfaces [ST91]. Intermediate [BV90b, Var87, Var90a, Var92]. Internals [GC94]. internationella [Hol88]. Internet [CBR03, Keh93]. Interpretation [AwJS85, MB90, dB90]. Interpreter [Kam90]. Interpreter-Based [Kam90]. Introduction [Bea88, Big89, Bud91, CGG^{+}92, CW83, CLR89, Cra89, Dev87, Dev89, GO92, GT88b, Hec93, Hel90, Knu08, McC90, NL90, Sok89, Spe94, WF82, Wei91, Wig90, Dou90, Pif91]. Introduzione [RFB91]. invariant [Bea88]. Investigations [Cuo89, Smi90]. Ion [BW91]. IR [Bau89]. Ischaemic [C12te]. ISO [App90]. ISO-Norm [App90]. Issues [Cal88]. istorii [GKL^{+}92]. itinerario [Boi90].

jag [Ber89]. jej [Bie91]. Jewr’ias [LP92]. Joint [Alf92]. Journal [C65te, C76te, JCL89, Wol90b, UMA, JCo92, Sol92]. Journey [Gud94]. Juan [Est90]. June [Wel89a].

Kak [Sos92]. KGB [Rau92]. Kinematics [McC90]. Knowledge [Bac90, Ber85, BLR91, CRR^{+}90, DL82, PSF91]. Knowledge-Based [CRR^{+}90, DL82]. komputerowej [Bie91]. Koncepcja [Bie91]. Konstnärinna [Ber89]. krepostnojj [Pet92a]. kulqture [GKL^{+}92]. kul’tury [L^{+}92]. Kurt [FDK^{+}86].

Laboratorio [Boi90]. Laboratory [New88]. Lagrangian [Dou90]. Lambeth [WAD16]. LAN [For89]. Lanalisi [LN91]. Landscapes [PK91]. Language [BBR87, BW84, Car90, CGG^{+}91b, Dyb87, GFOE87, Hil94, JCL89, Ste90a, WWF87, Sos92, Kah98]. Languages [BB89, BS91, Kam90, LMW88, Syz91]. LAPACK [ABB^{+}92]. Laser [Kru88]. Lasers [Alf92]. Lattice [DFH^{+}90]. Laughing [LL87]. Laughter [Str92]. Learning [Jud90]. Leaves [CGG^{+}92]. Lecture [Dev86, MW91]. Lectures [C77te, Jen90, Van89]. Ledgers [ML91]. Legal [Ash90]. Legged [Rai86]. Leistungsbewertung [wHR89]. Lenguaje [W^{+}86]. Lenin [Vol92]. Leonardo [Sig02]. lexicographica [Mil96]. Liber [Sig02]. Library [CGG^{+}91a, Pal89, Win89a, wJBH86, WAD16]. Lie [BdK90, SMW86, Wol89]. Life [Lan89]. Light [His96]. Lineamenti [OP88]. Linear [Ber91, Dev90, DDSv91, GMW91, NM91]. Linguistic [BW84]. Linguistics [All91]. Linux [JT98, Stu01]. Lipid [C64te]. Lipidology [C49te]. Liquids [HLZJ91]. Lisp [Bro85, WH89, WH88, Ste90a]. LISPer [FF87]. Little [FF87]. Logic [Ric89, RBM^{+}91, Tic91]. Logical [Hue89, MW90a, MW90b]. Logo [Har85, Har86, Har87]. Low [RS88].

Mach [BKLL93]. Machine

[AK91, Hil85, MB90]. Macromolecular [C7890]. magazine [TVG87]. Magic [GC94]. MAKE [TNY92, TNY94]. Management [C64te, Lea85]. Manifolds [Mun91]. Manipulation [MS85]. Manipulator [AAH88]. Manipulators [Pau81]. Manual [CGG^{+}91a, CGG^{+}91b, C25te, C26te, C27te, C28te, C29te, C30te, C31te, C32te, C33te, KN94, SS89, W^{+}86, wJBH86]. manuals [VMS88]. Manuscript [WAD16]. Manuscripts [Pal89, Win89a]. Many [NO88, SJ87]. Many-Body [SJ87]. Many-Particle [NO88]. Maple [Hec93, KN94, GH93, CGG^{+}91a, CGG^{+}91b, CGG^{+}92]. Mark [Ahl88]. Marshall [Van89]. Marwan [Con89]. Marx [And90]. Mastering [Jae89, TNY92, TNY94]. Matematica [Pav92, Boi90]. matematicheskuyu [Sos92]. Mathematica [Mae90, Wol90b, Wol88, Wol90a, Rob95, Var93]. Mathematicae [CMU89]. Mathematical [AMS8x, Gud94, Kra01, LS89, MG89, OALD97, Pav92, vN18, Pav93, Sos92]. Mathematics [DW90b, GKP89, GKP94, GK82, GK81, Hoe90, Kee88, KN94, MR91, Pau81, Pif91, Wat90, Wei99, Wol88, Wol90a, WL89, Sol92, Rob95]. Mathematik [LHR^{+}91a, LHR^{+}91b]. MATLAB [GH93, SB90]. Matrix [GV89]. Matter [LL87, RS88, Tes90]. Measurement [Hil84]. meccanici [CV90]. Mechanics [Cra89, MS85, Str89, dB90, vN18]. Mechanik [SS89]. Mechanik-Manual [SS89]. Media [HR90]. Medical [Vol92]. Medicine [C62te, Pal89, Wel89b, Wel89a, Win89a, wJBH86]. meditsinskii [Vol92]. Memory [DDSv91, ES90]. Methodology [WWF87]. Methods [BA96, BBBH92, Cal88, DFH^{+}90, GO92, GT88b, Gra90a, LS89, SMW86, Wei90a, Wol89, Sal01]. MICROCALC [Boi90]. Microcomputer [CDF89, Wak89]. Microeconomia [Var90b, BV90a]. Microeconomic [Kre90, Var91]. Microeconomics [BV90b, Var87, Var90a, Var92]. microfiche [ML91]. Microwave [Poz90]. mid [Stu16].

Migraine [C30te]. Millennium [Knu05a]. millisecond [BvdH91]. Mirror [Wel89b, Wel89a]. MMIX [Knu05a]. Mobile [Aya91]. Model [AAH88, Sab88, RR89]. Model-Based [AAH88]. Modeling [Ash90, Bar90, Kah98, MC88, MG89, OALD97, Sal99, Var93]. Modelli [CV90]. Models [Bha88, BP89b, CH93, HR90]. Modern [BA96, Con89, Hel90, Kan88, Ram90, Sig02, Var92]. Moderno [Bau89]. Modularization [Gri89]. Molecular [PK91]. Molekulare [LB91]. Monday [Wel89b, Wel89a]. mondo [FR89]. Monitoring [C05te]. morfologicznej [Bie91]. morphological [Bie91]. Motion [BHJ^{+}82, Hil84, MB90]. MP/Model [RR89]. MS [JC88]. MS-DOS [JC88]. Multivariate [Wei90a]. Musical [Ber93]. Myocardial [C25te].

N [Van89]. napisat [Sos92]. National [Prz91, Gra88]. nationalekonom [And90]. NATO [BB90, BBBH92]. Natural [BBR87, Car90, GFOE87]. Network [Ber91, Jud90]. Networks [FS91, Gir90, Pao89]. Neural [FS91, HPK91, Mea89, Pao89, Jud90, NC89]. Neurobiology [C50te]. Neurocomputing [HN90]. Neurology [C51te]. Neurosurgery [C51te]. No [LL87]. Non [NM91]. Non-Linear [NM91]. Nonlinear [Wig90, Wol83]. Norm [App90]. North [GKL^{+}92]. Notes [Dev86, MW91]. novel [Rau92]. Nuclei [SJ87]. Nucleon [Bha88]. Numerical [GMW91, GO92, Gra90a, Hig96, Par93, PC89, PFTV86, Sad92, Thi88]. Nutritional [Gib90].

Oberflächenphysik [HwCZ91]. Object [Bud91, Gri91, Kir89]. Object-Oriented [Bud91, Kir89]. Objektorientierte [Kah98]. Obstetrics [C52te]. Oceanographic [DTZJ90]. October [Wel89b]. ODA [App91a, App90, App91b]. Offenen [App90]. Office [App90, App91b]. Old [Ste92]. Oncology [C53te]. Online [C61te]. Ontic [McA89]. Open [App91a, GC94, JA90, vMR90]. Operator [Dev90, Bea88]. Ophthalmology [C54te, C8390, C8491]. Opinion [C40te, C41te, C42te, C43te, C44te, C45te, C46te, C47te, C48te, C49te, C50te, C51te, C52te, C53te, C54te, C55te, C56te, C57te, C58te, C59te, C60te]. Optics [NM91, SMW86, Wol89]. Optimal [ACC^{+}95]. Optimization [Ber91, GMW91, GMW81, Nak91, Par93]. Oral [Stu16]. ordinary [HW91]. Organization [Wie87]. Oriented [Bud91, Kir89]. origins [Gra88]. Orthopaedics [C55te]. OS-9 [Dib92]. OS/2 [TNY94]. Osteoarthritis [C31te]. other [Win89a]. our [FvGGM90]. outils [ASU89]. own [Stu16].

Packaging [Bak90]. Paintings [Win89a]. Palace [WAD16]. Papers [C17te, C18te, C19te, C20te, Knu00, Sos92]. Paraguay [Nic93]. Paralation [Sab88]. Parallel [CG90, CM88, FWM94, KS91, SBK89, Syz91, Tic91]. Part [BdK90, GT88b]. Partial [DFH^{+}90]. Particle [Kan88, NO88, PDG90]. Particles [Hat91, Van89]. Partitions [Knu05b]. Parts [BW91, HLZJ91]. Pascal [Kay85, Kel82, Big89]. passivi [Bec88]. Patent [C10te, C09te, C11te, C15te, C34te, C16te, C70te, C79te]. Patents [C61te, C63te]. Paths [Tho88]. Pattern [Pao89]. pc [Cla86]. pc-Portable [Cla86]. Pearl [Smi90]. Peasant [Cas90]. Pediatrics [C56te]. Peninsular [Bau89]. People [Kay85]. People/Pascal [Kay85]. Peptic [C32te]. Perception [Con89]. Performance [BW84, KS91]. Personal [Wat90, Pic92, Stu16]. Perspective [WB79]. perspectives [Cla84]. pertinentia [Mil96].

Perturbation [BA96]. Perturbative [Fie89]. Pesticide [C79te]. Phenomena [Wol83]. Phoenix [Dai90]. Phonological [Lim10]. Phonology [BS91, Lim10]. Physical [GT88b]. Physically [Bar90]. Physically-Based [Bar90]. Physics [Alf92, ET90, ET91, Kan88, KM90, Kru88, MC91, RS88, SJ87, Sok89, Taj89, Tes90, Zur90, BdK90]. Physik [SKPS91]. Physiology [OALD97]. Pisano [Sig02]. Plan [Car90]. Planning [BHJ^{+}82]. Plant [C70te]. Plasma [Kru88, Taj89]. Plasmas [Par91, Van89]. PLUS [Mat85, Spe94]. po-anglijski [Sos92]. Pocket [C80te]. Point [Hat91]. Policies [ACC^{+}95]. Polish [Bie91]. political [Vol92]. politicheskii [Vol92]. Portable [Cla86]. Power [KSV91]. Practical [GMW81, PC89, van90, vH94]. Practice [C36te, C62te, Ste90b, wJBH86]. Practised [wJBH86]. Precalculus [DW90a, DW90b]. predictive [Bie91]. prepare [Sos92]. Pressure [C05te, C68te]. Prichernomorqja [GKL^{+}92]. Prilozhenie [L^{+}92]. Primer [Ram90]. Principes [ASU89]. Principle [dB90]. Principles [ASU89, Gib90, KSV91, Kee88]. printers [HH93]. printing [ML91, PJS93]. Privacy [Sal03]. Probabilistic [Bac90, BBBH92, dB90]. Probability [Ham91, Gud88]. Problems [GH93, vMR90, HW91, Pav93]. Procedural [Rog85]. procedures [L^{+}92]. Proceedings [BCH^{+}90, BB90, BBBH92]. Processing [GFOE87, Sch91, App91b]. progetto [Bie87]. Program [Cal88, CM88, KR99]. Programmer [RW90]. Programmers [Dib92]. Programming [BB89, BKLL93, Bro85, Bud91, Dyb87, Gra90a, Hil94, Hue89, Kam90, Kay85, Kel82, Kir89, Knu05a, Knu05b, Knu08, Knu09, LMW88, Mae90, MW90a, MW90b, Pau81, SF89, Syz91, Tho91, Tic91, Tur90, Wak89]. Programowanie [WG93]. Programs [AwJS85, CG90, Dij89, TNY92, TNY94].

Progress [HvM92]. Projections [Sal06b]. Projects [Gra90a]. Prolog [O’K90, SS86, Ste90b]. Proofs [Dij89]. Propagation [Bir91]. Properties [PDG90]. Proteins [PK91]. Protocols [JA90]. Prozeßautomatisierung [BV91]. przykładach [WG93]. Psychiatry [C57te]. Publications [FDK^{+}86]. Publishers [Fly95, HH93]. pulsars [BvdH91].

QCD [Fie89]. Quantum [Gud88, Hat91, NO88, vN18]. Quarks [Bha88]. Queuing [Dai91].

R [Con89]. radio [BvdH91]. Radiology [C58te, C8590, C8691]. Rasprostranenie [Bir91]. Ray [Sok89]. RC [Bie87]. RC-attivi [Bie87]. RÉ [Bau89]. Reactions [BW91]. readings [Ram98]. Reaktionsdynamik [LB91]. Real [Tor88]. Reasoning [Bac90]. Rechensystemen [wHR89]. Recipes [PFTV86]. Recognition [Car90, Gri91, Pao89]. recollections [Stu16]. REDUCE [MW91]. Redundancy [Nak91]. Reference [CGG^{+}91a, CGG^{+}91b, Sal98, Sal00, Sal04, Sal07]. Reflections [ES90]. register [HH93]. Regulator [C70te]. relativistic [Dou90]. Release [GC94]. Repelling [CBR03]. Representation [BLR91]. Representing [Bac90]. reproduced [ML91]. Research [Tur90, Cla84]. Reserve [ACC^{+}95]. resources [Wel89b, Wel89a]. Restructuring [Lim10]. Result [AK93]. results [Kup87]. Retargetable [FH95]. reti [BB88]. Retrospect [Gra88]. Reusability [BP89a, BP89b, Gri89]. Review [She92]. Rheumatoid [C33te]. Rheumatology [C20te, C24te, C59te]. RISC [Knu05a]. Risk [Cas90, C36te]. Rituals [Les91]. RNA [PK91]. Robot [AAH88, BHJ^{+}82, HS91, Hor86, MS85, Pau81]. Robotics [Cra89, Nak91]. Robots [Aya91, Rai86]. Role [Gri91]. Roles [Les91]. Romanum [Mil96]. Rosenbluth [Van89]. Routing [Gir90]. rozproszone [WG93]. Rugged [PK91]. Rulfo [Est90]. Russian [Vol92]. Russians [Pet92b]. Russkie [Pet92b].

S [CH93, Mat85, Spe94]. S-PLUS [Mat85, Spe94]. salute [FvGGM90]. Scheme [SF89, Dyb87]. School [MW91]. Science [GT88a, Har85, Har86, Har87, PJS92, PJS93, Ras91]. Sciences [Kra01]. Scientific [AK93, GH93, GO92, Nas90, PFTV86]. Scientists [BW90, Ham91, Pie91]. Sea [GKL^{+}92]. Search [Aig88]. Searching [Knu98b]. second [PJS93]. Security [CBR03, Sal03, JCo92, Sal05b]. segnali [LN91]. sekelskiftets [And90]. Selected [Knu00]. Semiconductor [Alf92]. Seminumerical [Knu98a]. September [Wel89a]. Serf [Pet92a]. Series [New88, Sch91, Wei90a, C39te, Stu16]. Service [Gra88]. Seventh [KN94]. Severnogo [GKL^{+}92]. SGML [van90, vH94]. Shading [HB89]. Shape [HB89, Koe89]. Shared [DDSv91]. Signal [Sch91]. Signals [Sie86]. signed [L^{+}92]. Simple [Ros91]. Simula [Kir89]. Simulation [CDF89, GT88b]. Simulations [ACC^{+}95]. Single [Ada89, Ros90]. Single-handedly [Ros90]. Sinification [Lim10]. Sinitic [Lim10]. Sintesi [Bec88]. sistemam [L^{+}92]. sistemi [CV90]. Slavic [ABS88]. Slide [C8390, C8491, C8590, C8691, C8190, C8291]. słownikowej [Bie91]. Social [MS88]. Society [AMS8x]. Software [BP89a, BP89b, Jon90, Kah98, WL89, Pif91, Kah98]. Software-Engineering. [Kah98]. Software-Entwicklung [Kah98]. Solid [Koe89]. Solitons [Bha88]. Solstice [Sol92]. Solved [Pav93]. Solving [DDSv91, GH93, HW91]. Some [Wel89b, Wel89a]. son [ML91]. Sonidos [Est90]. Sorting [Knu98b]. Soviet [Alf92].

Soviet-American [Alf92]. Space [Par91]. Special [BLR91]. Spectral [Bir91]. Spectral’naja [Bir91]. Speech [O’S87]. Spreadsheet [MC91]. SQL [Lus88]. Stability [Hig96]. Standard [App91a, Jae91, Jae89]. Stat’i [L^{+}92]. Statistical [CH93, HR90, Sch91, Thi88]. Statistics [Wei91]. stat’yu [Sos92]. Steps [Sha90]. Stiff [HW91]. Stochastic [BBBH92, CW83]. Stoiqe’ia [LP92]. Stories [Dai92, Str92]. Strata [MS88]. Stratification [MS88]. Strings [Hat91]. Strong [SJ87]. Strongly [Fra91]. Structural [C60te]. Structure [AwJS85]. Structures [C7890, Dal87, WH90, KR99]. Studies [BS91, Con89, GKL^{+}92, OALD97, ABS88]. Study [BB90, BBBH92, Gri81]. Style [Har85, Har86, Har87]. subspaces [Bea88]. Superconductivity [BCH^{+}90, Fra91, Hal88, OD91]. superintegrable [Kup87]. Supplement [Pal89]. Supplementary [Pal89]. Support [CDF89]. Surfaces [Gri81, Sal06a]. Surgery [C0790]. Survey [GP87]. svenska [Hol88]. svolti [Pav93, SV91]. Switched [Gir90]. symmetry [Dou90]. Syntactic [Ber85]. System [AAP88, GC94, Gri81, PK91, Wol88, Wol90a, Ram98]. Systematic [Jon90]. Systemen [App90]. Systems [App91a, CRR^{+}90, DL82, Dev87, Dev89, DDSv91, Fra91, FPW90, FPEN91, GC94, GT88b, HWH91, JA90, Jen90, KS91, Mea89, NO88, Omo94, PC89, Sie86, SBK89, TNY92, TNY94, WF82, Wig90, App91b, Kup87, L^{+}92, CS87].

Tables [AAdR85]. Taiwan [Dai92]. Tears [Str92]. technical [Ram98]. Techniques [Knu09, RS88, Sad92, Stu01, Syz91, ASU89, Kup87, Pic92]. Technologies [Prz91]. Technology [Prz91, For89, Rah87]. Tecniche [Pic92]. Temi [Pav93, SV91]. Temperature [BCH^{+}90, Fra91, Hal88].

Temperatures [RS88]. Tentative

[AAdR85]. Teoria [Bie87, BB88]. teorija [Bir91]. TEXbook [Sal95]. Text [App91b]. Texts [Knu91, She92]. th [Lim10]. Thankas [Win89a]. their [Stu16]. Theoretical [ET90, ET91, McC90, Nak91, Per88a, Per88b]. Theories [BW85, Hal88, Tes90, Wel90]. Theory [ACC^{+}95, BW90, Ber93, Bir91, Dai91, Dev90, Gut90, Hat91, Kre90, Lea85, LP92, Pie91, Ram90, Tho91, Bea88]. Therapeutic [C63te]. Thought [SHH^{+}89]. Tibetan [ES90, Win89a]. Time [GKL^{+}92, New88, Sch91, Wei90a]. Timeslab [New88]. Tipografía [W^{+}86]. tipologii [L^{+}92]. Tips [Stu01]. Tomography [DTZJ90]. Tools [JC88, ASU89]. Topics [Bar90, Tur90]. Topology [HvM92, vMR90]. Toposes [BW85]. Tra [FR89]. Transactions [IEE89]. Transform [Bra86, Van92]. Transformations [Sal06b]. Transition [HLZJ91]. Translation [Cal88, Sig02, JCL89]. Tree [Dai90]. Trends [ET90, ET91]. Tribal [Cas90]. Tricks [Knu09]. Trigonometry [DW89b, DW89c]. Triples [BW85]. Trudam [L^{+}92]. Tubes [Gra90b]. Turkic [Lim10]. Turrell [ES90]. Turtle [Ad81]. Tutorial [CGG^{+}92]. TV [TVG87]. ty [Abr91]. Type [Tho91, Moy95]. types [L^{+}92]. Typography [Kra01].

UK [Cla84]. Ulcer [C32te]. Ultrahigh [Sok89]. UMAP [UMA]. Uncertainty [Cas90, dB90]. Unified [Kah98]. Univariate [Wei90a]. Universe [KT90]. Universitatis [CMU89]. UNIX [AL91, GC94, TNY92, TNY94]. unofficial [Seh92]. Unumane [Ros90]. uravnenija [Bir91]. Use [BW84, Stu01]. User [ABB^{+}92, Mat85]. Users [Fly95]. Uses [WP84]. Using [GH93, Jon90, Kel82, Pic92]. uso [Pic92]. Utrikespolitk [Hol88]. v [Pet92b, Rau92, CGG^{+}91a, CGG^{+}91b, CGG^{+}92, ES90, GC94]. V. [Vol92]. Variability [C05te]. Variable [Ada89]. Variables [Tor88]. Vävarfolket [Ahl88]. VAX [JC88, VMS88]. VAX/VMS [JC88]. VDM [Jon90]. Vector [DDSv91]. Vegas [Seh92]. Vek [Pet92b]. Veka [Pet92a]. vem [Ber89]. Vent [Ham90]. Vento [TP90a]. Verification [AK93, Bie91]. Version [Mat85, VMS88]. Vestnik [EVxx]. vet [Ber89]. villagers [Stu16]. Vision [Aya91, Cha91, HS91, Hor86]. Visual [Gri81, Hil84, MB90]. Visualization [KS91]. VLSI [Bak90, Dal87, Mea89]. VMS [JC88, VMS88]. Vol [FDK^{+}86, Wol89]. voln [Bir91]. Volume [Knu05a, Knu05b, Knu08, Knu09, BLR91]. volumes [WB79]. Vremya [Abr91]. vspyat [Abr91].

W [FvGGM90, WG93]. Warren [AK91]. Wave [Bir91, dB90]. Wellcome [Pal89, Wel89b, Wel89a, Win89a, wJBH86]. weryfikacji [Bie91]. Western [Pal89]. Wien [And90]. William [ML91]. Wily [CBR03]. Windatlas [TP90c]. Windows [Mat85]. Women [Les91]. words [Stu16]. Workouts [BV90b]. Works [FDK^{+}86, FWM94]. Workshop [Alf92, Tes90]. World [Con89, FR89]. WorldWideWeb [Fly95]. Write [CG90]. Written [Lim10]. współbieżne [WG93]. Wylie [ES90].

X [RR89]. X-MP [RR89]. X-MP/Model [RR89]. Xylographs [Win89a].

York [HH93]. Ypologismo’u [LP92].

zadaniach [WG93]. Zen [Keh93]. znakovym [L^{+}92].

References

Angus:CTT85

[AAdR85] S. Angus, B. Armstrong, and K. M. de Reuck. Chlorine Tentative Tables. IUPAC chemical data series. Pergamon Press, Oxford, UK, 1985. ISBN 0-08-030713-2. xviii + 162 pp. LCCN QD181.C5.A54 1985. Typeset using TEX80.

An:MBC88

[AAH88] C. An, C. G. Atkeson, and J. M. Hollerbach. Model-Based Control of a Robot Manipulator. The MIT Press, Cambridge, MA, 1988. ISBN 0-262-01102-6. 233 pp. LCCN TJ211 .A551 1988. Prepared with TEX.

Anderson:EEC88

[AAP88] Philip W. Anderson, Kenneth J. Arrow, and David Pines. The Economy as an Evolving Complex System. Addison-Wes-ley, Reading, MA, USA, 1988. ISBN 0-201-15681-4 (hardcover), 0-201-15685-7 (paperback). xvii + 317 pp. LCCN HB21 .E2651 1988. Prepared with TEX.

Anderson:LUG92

[ABB^{+}92] E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, S. Ostrouchov, and D. Sorensen. LA-PACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-294-7. xv + 235 pp. LCCN QA76.73.F25 L36 1992. Prepared with L^{A}TEX.

Abragam:VVF91

[Abr91] A. Abragam. Vremya vspyat’, ili fizik, fizik, gde ty byl. Nauka, Glavnaya redakciya fiziko-matematicheskoj literatury, Moscow, Russia, 1991. ISBN 5-02-014712-5. 391 pp. Prepared with L^{A}TEX.Translated by the author from the original French edition, De la physique avant tout chose, Editions Odile Jakob.

ABSEES

[ABS88] The American Bibliography of Slavic and East European Studies. Library of Congress, 1988. ISSN 0094-3770. L^{A}TEX for text; plain TEX for 3-column index.

Alston:ORE95

[ACC^{+}95] J. A. Alston, H. F. Carman, J. E.

Christian, J. Dorfman, J.-R. Murua, and R. J. Sexton. Optimal Reserve and Export Policies for the California Almond Industry: Theory, Econometrics and Simulations. Giannini Foundation of Agricultural Economics, University of California, Agricultural and Natural Resource Publications, 6701 San Pablo Avenue, Oakland, CA 94608, USA, 1995. ISSN 0575-4208. LCCN HD1407 .G5. US$14.00.

Abelson:TG81

[Ad81] Harold Abelson and Andrea A.

diSessa. Turtle Geometry. The MIT Press, Cambridge, MA, 1981. ISBN 0-262-01063-1. xx + 477 pp. LCCN QA462 .A23 1981. This book was prepared with TEX80; this was described in TUGboat 2, no. 3 in an article by Michael Sannella.

Adams:SVC89

[Ada89] Robert Adams. Single Variable Calculus. Addison-Wesley, Don Mills, Ontario, Canada, 1989. ISBN 0-201-11114-4. Prepared with TEX.

Ahlberger:1988:VVF

[Ahl88] Christer Ahlberger. Vävarfolket: Hemindustrin i Mark 1790– 1850. Number 1 in Skrifter från Institutet förr lokalhistorisk forskning. De sju häradernas kulturhistoriska förening, Gothenburg, Sweden, 1988. ISBN 91-7900-417-2. 199 pp. Distribution: Borås Museum, Ramnaparken, S–502 65 Borås, Sweden.

Aigner:CS88

[Aig88] Martin Aigner. Combinatorial Search. John Wiley and B. G. Teubner, New York, NY, USA and Stuttgart, Germany, 1988. ISBN 0-471-92142-4 (Wiley), 3-519-02109-9 (Teubner). 368 pp. LCCN T57.97 .A381 1988. Prepared with L^{A}TEX.

Hassan:WAM91

[AK91] Hassan Ait-Kaci. Warren’s Abstract Machine. The MIT Press, Cambridge, MA, 1991. ISBN 0-262-01123-9 (hardcover), 0-262-69146-9 (paperback). xvi + 114 pp. LCCN QA76.73.P76 A38 1991. Prepared with L^{A}TEX.

Adams:1993:SCA

[AK93] E. Adams and U. Kulisch. Scientific Computing with Automatic Result Verification, volume 189 of Mathematics in Science and Engineering. Academic Press, Boston, MA, USA, 1993. ISBN 0-12-044210-8. x + 612 pp. LCCN QA76 .S368 1993. Edited by William F. Ames.

Abrahams:UI91

[AL91] Paul W. Abrahams and Bruce R.

Larson. UNIX for the Impatient. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-55703-7. xxvii + 559 pp. LCCN QA76.76.O63 A27 1992. Prepared with TEX.

Alferov:JSA92

[Alf92]  Zhores I. Alferov.  Joint

Soviet-American Workshop on the Physics of Semiconductor Lasers. American Institute of Physics, Woodbury, NY, USA, 1992. ISBN 0-88318-936-4. viii + 183 pp. LCCN TA1700 .J65 1991.

Computational-Linguistics

[All91] James S. Allen. Computational linguistics, 1991. ISSN 0362-613X. To be published quarterly by MIT Press; fonts not yet finalized.

AMS

[AMS8x] American Mathematical Society, 198x. By the end of the 1980s, almost all AMS journals and monographics have been set using TEX. This entry needs to be replaced by an explicit list of them.

Andren:BME90

[And90] Mats Andr’en. Borgerskapets Marx—Eugen von Böhm-Ba-werk: ämbetsman och nationalekonom i sekelskiftets Wien. Symposion Bokförlag, Stockholm/Stehag, Sweden, 1990. ISBN 91-7868-226-6. English title: The Bourgeois Marx— Eugen von Böhm Bawerk: Official and Economist in the Late Habsburg Vienna.

Angel:CG89

[Ang89] Edward Angel. Computer Graphics. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-13548-5. xviii + 477 pp. LCCN T385.A513 1990. Prepared with L^{A}TEX.

Appelt:DOS90

[App90] Wolfgang Appelt. Dokumentaustausch in Offenen Systemen —Einführung in die ISO-Norm 8613: Office Document Architecture (ODA) and Interchange Format. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 3-540-52707-9 and 0-387-52707-9. Prepared with TEX.

Appelt:DAO91

[App91a] Wolfgang Appelt. Document Architecture in Open Systems: The ODA Standard. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-54539-5 and 0-387-54539-5. x + 350 pp. LCCN TK5105 .A66 1991. Prepared with TEX.

Appelt:ISO-IEC-8613

[App91b] Wolfgang Appelt. Information processing—text and office systems—office document architecture (ODA) and interchange format. Part 10: Formal specifications Standard 8613, International Organization for Standardization, Geneva, Switzerland, 1991. Prepared with TEX.

Ashley:MLA90

[Ash90] Kevin Ashley. Modeling Legal Argument. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-01114-X. x + 329 pp. LCCN KF242.A1 A7 1990.

Aho:CPT89

[ASU89] Alfred Aho, Ravi Sethi, and Jeffrey Ullman. Compilateurs—Principes, techniques et outils (English: Compilers —Principles, Techniques and Tools). InterEditions, 25 rue Leblanc, 75015 Paris, France, 1989. ISBN 2-7296-0295-X. 875 pp. Prepared with TEX.Translated from the English edition (Addison-Wesley, 1986, ISBN 0-201-10088-6) by Pierre Boullier, Philippe Deschamp (sic), Martin Jourdan, Bernard Lorho and Monique Mazaud.

Astrom:AC89

[ÅW89] Karl Johan Åström and Björn Wittenmark. Adaptive Control. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-09720-6. xiv + 526 pp. LCCN TJ217 .A671 1989. Prepared with TEX.

Abelson:SIC85

[AwJS85] Harold Abelson and Gerald J. Sussman with Julie Sussman. Structure and Interpretation of Computer Programs. The MIT Press and McGraw-Hill, Cambridge, MA, 1985. ISBN 0-262-01077-1 (MIT Press), 0-07-000422-6 (McGraw-Hill). xx + 542 pp. LCCN QA76.6 .A255 1984. This book was prepared with TEX80.

Ayache:AVM91

[Aya91] Nicholas Ayache. Artificial Vision for Mobile Robots. The MIT Press, Cambridge, MA, 1991. ISBN 0-262-01124-7. 342 pp. LCCN TJ211.3 .A9313 1991. Prepared with L^{A}TEX.

Bond:1996:MAF

[BA96] V. R. Bond and M. C. Allman. Modern Astrodynamics: Fundamentals and Perturbation Methods. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-04459-7. xi + 250 pp. LCCN TL1050.B66 1996. US$35.00,

                                        UK£27.50. URL http://aaup.
                                        pupress.princeton.edu:70/
                                        0h/books/presses/princeton/
                                        titles/95031024.html. Pre-

pared with L^{A}TEX.

Bacchus:RRP90

[Bac90] Fahiem Bacchus. Representing and Reasoning with Probabilistic Knowledge. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-02317-2. 233 pp. LCCN QA273 .B24 1990. Prepared with L^{A}TEX.

Bakoglu:CIP90

[Bak90] H. B. Bakoglu. Circuits, Interconnections, and Packaging for VLSI. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-06008-6. xiv + 527 pp. LCCN TK7874.B345 1990. Prepared with L^{A}TEX.

Barr:TPB90

[Bar90] Alan H. Barr, editor. Topics in Physically-Based Modeling. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-51013-8. Prepared with L^{A}TEX.

Bauhr:FRE89

[Bau89]  Gerhard Bauhr.  El Futuro

en -RÉ e IR A + Infinitivo en Español Peninsular Moderno. Number 39 in Romanica Gothoburgensia. Acta Universitatis Gothoburgensis, Gothenburg, Sweden, 1989. ISBN 91-7346-218-7. ISSN 0080-3863. ix + 405 pp. LCCN PC 4271 B38 1989.

Beccari:ETR88

[BB88] C. Beccari and M. Biey. Esercitazioni di teoria delle reti elettriche. Edizioni Cooperativa Libraria Universitaria Torino, Corso Duca degli Abruzzi, 24, I 10129 Torino, Italy, 1988. Prepared with L^{A}TEX.

Bancilhon:ADP89

[BB89] François Bancilhon and Peter Buneman, editors. Advances in Database Programming Languages. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-50257-7. xvii + 457 pp. LCCN QA76.9.D3 A348 1989. Prepared with TEX.

Byrnes:Fourier-ASI89

[BB90] J. S. Byrnes and Jennifer L. Byrnes, editors. Recent Advances in Fourier Analysis and its Applications: Proceedings of the NATO Advanced Study Institute, volume 315 of NATO ASI series C: Mathematical and physical sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, Il Ciocco Resort, Tuscany, Italy, 1990. ISBN 0-7923-0875-1. LCCN QA403.5 .N37 1989.

Byrnes:Prob-Sto-ASI91

[BBBH92] J. S. Byrnes, Jennifer L. Byrnes, Karl Berry, and Kathryn A. Hargreaves, editors. Probabilistic and Stochastic Methods in Analysis, with Applications: Proceedings of the NATO Advanced Study Institute, volume 372 of NATO ASI series C: Mathematical and physical sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, Il Ciocco Resort, Tuscany, Italy, 1992. ISBN 0-7923-1804-8. LCCN QA274.2 .P77 1992.

Berwick:CCN87

[BBR87] R. C. Berwick, G. E. Barton, and E. S. Ristad. Computational Complexity and Natural Language. The MIT Press, Cambridge, MA, 1987. ISBN 0-262-02266-4. xii + 335 pp. LCCN QA267.B36 1987. Prepared with L^{A}TEX.

Bedell:HTS90

[BCH^{+}90] K. S. Bedell, D. Coffey, D. E.

Heltzer, D. Pines, and J. R. Schrieffer. High Temperature Superconductivity Proceedings. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-51249-1. xiv + 474 pp. LCCN QC611.98.H54 H34 1989. Prepared with TEX.

Bauerle:LA190

[BdK90] G. G. A. Bäuerle and E. A.

de Kerf. Lie Algebras Part 1 —Finite and infinite dimension Lie algebras and applications in physics. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1990. ISBN 0-444-88776-8. xvi + 394 pp. LCCN QC20.7.L54 B38 1990. Studies in Mathematical Physics, volume 1.

Beauzamy:IOT88

[Bea88] Bernard Beauzamy. Introduction to operator theory and invariant subspaces, volume 42 of North-Holland Mathematical Library. North-Holland Publishing Company, Amsterdam, The Netherlands, 1988. ISBN 0-444-70521-X. xiv + 358 pp. LCCN QA329 .B441 1988. Prepared with TEX.

Beccari:SCP88

[Bec88] C. Beccari. Sintesi dei circuiti passivi. Edizioni Cooperativa Libraria Universitaria Torino, Corso Duca degli Abruzzi, 24, I 10129 Torino, Italy, 1988. Prepared with TEX by Tecnograf.

Berwick:ASK85

[Ber85] Robert C. Berwick. The Acquisition of Syntactic Knowledge. The MIT Press, Cambridge, MA, 1985. ISBN 0-262-02226-5. xii + 368 pp. LCCN Q335 .B48 1985. Prepared with TEX.

Bergstrom:EFJ89

[Ber89] Irja Bergström. Ensam fågel— jag vet vem det är: Ester Henning Konstnärinna på hospital. Acta Universitatis Gothoburgensis, Gothenburg, Sweden, 1989. ISBN 91-7900-753-8. English title: Lonely bird—I know who that is. Doctoral dissertation. 140 illustrations, 20 in color.

Bertsekas:LNO91

[Ber91] Dimitri Bertsekas. Linear Network Optimization. The MIT Press, Cambridge, MA, 1991. ISBN 0-262-02334-2. xi + 359 pp. LCCN T57.85 .B43 1991. Prepared with L^{A}TEX.

Berman:MI93

[Ber93] Laurence Berman. The Musical Image: A Theory of Content. Contributions to the study of music and dance. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1993. ISBN 0-313-28434-2. ISSN 0193-9041. 391 pp. LCCN ML3845 .B377 1993.

Buckley:DG90

[BH90] Fred Buckley and Frank Harary.

Distance in Graphs. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-09591-2. xiii + 335 pp. LCCN QA166 .B83 1990. Prepared with TEX.

Bhaduri:MNF88

[Bha88] Rajat K. Bhaduri. Models of the Nucleon: From Quarks to Solitons. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-15673-3. xvii + 360 pp. LCCN QC793.2 .B48 1988. Prepared with TEX.

Brady:RMP82

[BHJ^{+}82] Michael Brady, John M. Hollerbach, Timothy L. Johnston, Tomas Lozano-Perez, and Matthew T. Mason, editors. Robot Motion: Planning and Control. The MIT Press, Cambridge, MA, 1982. ISBN 0-262-02182-X. xv + 585 pp. LCCN TJ211 .R54 1982. Prepared with TEX.

Biey:CRC87

[Bie87] M. Biey. Circuiti RC-attivi, Teoria e progetto. Edizioni Cooperativa Libraria Universitaria Torino, Corso Duca degli Abruzzi, 24, I 10129 Torino, Italy, 1987. Prepared with TEX by Tecnograf.

Bien:KSI91

[Bie91]  Janusz S. Bień.  Koncepcja

słownikowej informacji morfologicznej i jej komputerowej weryfikacji. (Polish) [The concept of predictive information and its morphological verification of computer], volume 383 of Rozprawy Uniwersytetu Warszawskiego. Wydawnictwa Uniwersytetu Warszawskiego, Warszawa, Polska, 1. edition, 1991. ISBN 83-230-0390-4. ISSN 0509-7177. składał: autor.

Biggs:ICP89

[Big89] Norman L. Biggs. Introduction to Computing with Pascal. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1989. ISBN 0-19-853755-7 (hardcover), 0-19-853756-5 (paperback). x + 219 pp. LCCN QA76.73.P2 B54 1989. Prepared with L^{A}TEX.

Birman:DU91

[Bir91] Differentsial’nye uravnenija. Spectral’naja teorija. Rasprostranenie voln. (English: Differential Equations. Spectral Theory. Wave Propagation), page 310, 1991. ISSN 0555-2818. Leningrad University, St. Petersburg, Russia.

Bruggeman-Klein:ED89

[BK89] Anne Brüggeman-Klein. Einführung in die Dokumentenverbeitung. B. G. Teubner, Stuttgart, Germany, 1989. ISBN 3-519-02488-8. 200 pp. Prepared with L^{A}TEX.

Boykin:PM93

[BKLL93] Joseph Boykin, David Kirschan, Alan Langerman, and Susan LoVerso. Programming under Mach. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-52739-1. xvii + 490 pp. LCCN QA76.8.N49 P76 1993.

Brachman:KR91

[BLR91] Ronald J. Brachman, Hector J.

Levesque, and Raymond Reiter, editors. Special volume: Knowledge Representation. Elsevier Science Publishers, Amsterdam, The Netherlands, 1991. ISBN 0-262-52168-7, 0-262-02344-X. LCCN Q387.K56 1992. Artificial Intelligence 49 (1991) 1-3.

Bell:CDN90

[BM90] George I. Bell and Thomas Marr.

Computers and DNA. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-51505-9 (hardcover), 0-201-51561-X (paperback). xvii + 276 pp. LCCN QP624 .I53 1988. Prepared with TEX.

Boieri:LMI90

[Boi90] P. Boieri. Laboratorio di matematica, Un itinerario attraverso la geometria analitica e l’analisi con MICROCALC. Loescher Editore, Via Vittorio Amedeo II, 18, Torino, Italy, 1990. Prepared with L^{A}TEX.

Biggerstaff:SRA89-2

[BP89a] Ted J. Biggerstaff and Alan J.

Perlis, editors. Software Reusability: Applications and Experience, volume 2. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-50018-3. LCCN QA76.76.R47 B543 1989. Prepared with TEX.

Biggerstaff:SRC89-1

[BP89b] Ted J. Biggerstaff and Alan J. Perlis, editors. Software Reusability: Concepts and Models, volume 1. Addison-Wes-ley, Reading, MA, USA, 1989. ISBN 0-201-08017-6. LCCN QA76.76.R47 B543 1989. Prepared with TEX.

Bracewell:HT86

[Bra86] Ronald N. Bracewell. The Hartley Transform. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1986. ISBN 0-19-503969-6. vii + 160 pp. LCCN QA403.5 .B731 1986.

Brooks:PCL85

[Bro85] Rodney A. Brooks. Programming in Common Lisp. John Wiley and Sons, Inc., New York, NY, USA, 1985. ISBN 0-471-81888-7. xv + 303 pp. LCCN QA76.73.L23 B76 1985.

Boltz:1991:SHP

[BS91] William G. Boltz and Michael C. Shapiro. Studies in the Historical Phonology of Asian Languages, volume 77 of Current Issues in Linguistic Theory. John Benjamins, Amsterdam, The Netherlands, 1991. ISBN 90-272-3574-0 (Europe), 1-55619-132-4 (USA). viii + 249 pp. LCCN P381.A75 S78 1991.

Budd:IOO91

[Bud91] Timothy Budd. An Introduction to Object-Oriented Programming. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-54709-0. xv + 399 pp. LCCN QA76.64 .B83 1991. Prepared with L^{A}TEX.

Bergstrom:EM90

[BV90a] Theodore C. Bergstrom and Hal R. Varian. Esercizi di microeconomia. Libreria Editrice Cafoscarina, Venice, Italy, second edition, 1990. ISBN 88-85613-28-4. Prepared with TEX.

Bergstrom:WM90

[BV90b] Theodore C. Bergstrom and Hal R. Varian. Workouts in Intermediate Microeconomics. W. W. Norton & Co., New York, NY, USA, second edition, 1990. ISBN 0-393-95926-0. 384 pp. LCCN HB 172 B48 1990. Prepared with TEX.

Bolch:P91

[BV91] Gunter Bolch and Martina Vollath. Prozeßautomatisierung. B. G. Teubner, Stuttgart, Germany, 1991. ISBN 3-519-02499-3. Prepared with L^{A}TEX.

Bhattacharya:PR203-1-203

[BvdH91] D. Bhattacharya and E. P. J.

van den Heuvel (eds). Formation and evolution of binary and millisecond radio pulsars. Physics Reports, 203(1 & 2), 1991.

Berwick:GBL84

[BW84] R. C. Berwick and A. Weinberg. The Grammatical Basis of Linguistic Performance: Language Use and Acquisition. The MIT Press, Cambridge, MA, 1984. ISBN 0-262-02192-7. xviii + 325 pp. LCCN P158 .B48 1984. Prepared with TEX.

Barr:TTT85

[BW85] Michael Barr and Charles F. Wells. Toposes, Triples, and Theories. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-96115-1 and 3-540-96115-1. xiii + 345 pp. LCCN QA169 .B346 1985. Originally prepared in TEX78, but converted by publisher to L^{A}TEX.

Barr:CTC90

[BW90] Michael Barr and Charles F. Wells. Category Theory for Computer Scientists. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-120486-6. xv + 432 pp. LCCN QA76.9.M35B37 1990. Prepared with L^{A}TEX.

Broglia:HIR91

[BW91] R. A. Broglia and Åge Winther. Heavy Ion Reactions, Parts 1 and 2. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-51392-7. LCCN QC794.8.H4 B76 1991.

CurSci:AIDS

[C01te] AIDS, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. Current Science publishes more than 80 medical periodicals using an inhouse markup language for which TEX is the back-end typesetting system. All submitted articles (whether on paper or on magnetic media) are rekeyed by clerical staff into the markup system.

CurSci:AIDS-90

[C02te] AIDS ’90, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:AIDS-Afri

[C03te] AIDS in Africa, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:AIDS-A-Guid-Clin-Coun

[C04te] AIDS: A Guide to Clinical Counselling, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Ambu-Moni-Bloo-Pres-Vari

[C05te] Ambulatory Monitoring and Blood Pressure Variability, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Anae-Frie

[C06te] Anaesthetist’s Friend, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Annu-Card-Surg-1990

[C0790] Annual of Cardiac Surgery 1990, 1990. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Annu-Gast-Endo-1990

[C0890] Annual of Gastrointestinal Endoscopy 1990, 1990. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Anti-Pate-Fast

[C09te] Antimicrobial Patent Fast-Alert, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Anti-Horm-Pate-Fast

[C10te] Anticancer and Hormone Patent Fast-Alert, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Anti-Pate-Fast-2

[C11te] Anti-inflammatory Patent Fast-Alert, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Athe-Athe-Isch-Hear-Dise

[C12te] Atheroma: Atherosclerosis in Ischaemic Heart Disease, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Atla-Coro-Hear-Dise

[C13te] Atlas of Coronary Heart Disease, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Atla-Hear-Dise

[C14te] Atlas of Heart Disease in the Aged, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Biot-Pate-Fast

[C15te] Biotechnology Patent Fast-Alert, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Card-Pate-Fast

[C16te] Cardiovascular Patent Fast-Alert, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clas-Pape-Asth

[C17te] Classic Papers in Asthma, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clas-Pape-Hype

[C18te] Classic Papers in Hyperlipidaemia, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clas-Pape-Hype-2

[C19te] Classic Papers in Hypertension, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Clas-Pape-Rheu

[C20te] Classic Papers in Rheumatology, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Illu-Dict-Card

[C21te] Clinician’s Illustrated Dictionary of Cardiology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Illu-Dict-Gast

[C22te] Clinician’s Illustrated Dictionary of Gastroenterology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Illu-Dict-Hype

[C23te] Clinician’s Illustrated Dictionary of Hypertension, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Illu-Dict-Rheu

[C24te] Clinician’s Illustrated Dictionary of Rheumatology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Manu-Angi-Myoc

[C25te] Clinician’s Manual on Angina and Myocardial Infarction, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Manu-Asth

[C26te] Clinician’s Manual on Asthma, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Manu-Hear-Fail

[C27te] Clinician’s Manual on Heart Failure, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Manu-Hype

[C28te] Clinician’s Manual on Hyperlipidaemia, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Manu-Hype-2

[C29te] Clinician’s Manual on Hypertension, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Manu-Migr

[C30te] Clinician’s Manual on Migraine, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Manu-Oste

[C31te] Clinician’s Manual on Osteoarthritis, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Manu-Pept-Ulce

[C32te] Clinician’s Manual on Peptic Ulcer, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Clin-Manu-Rheu-Arth

[C33te] Clinician’s Manual on Rheumatoid Arthritis, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:CNS-Pate-Fast

[C34te] CNS Patent Fast-Alert, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Coro-Arte-Dise

[C35te] Coronary Artery Disease, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Coro-Risk-Asse-Prac

[C36te] Coronary Risk Assessment in Practice, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Biol

[C37te] Current Biology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Card-Drug

[C38te] Current Cardiovascular Drugs, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Drug

[C39te] Current Drugs (series), 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Anae

[C40te] Current Opinion in Anaesthesiology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Biot

[C41te] Current Opinion in Biotechnology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Card

[C42te] Current Opinion in Cardiology, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Cell-Biol

[C43te] Current Opinion in Cell Biology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Dent

[C44te] Current Opinion in Dentistry, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Gast

[C45te] Current Opinion in Gastroenterology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Gene-Deve

[C46te] Current Opinion in Genetics and Development, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Immu

[C47te] Current Opinion in Immunology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Infe-Dise

[C48te] Current Opinion in Infectious Diseases, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Lipi

[C49te] Current Opinion in Lipidology, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Neur

[C50te] Current Opinion in Neurobiology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Neur-Neur

[C51te] Current Opinion in Neurology and Neurosurgery, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Obst-Gyne

[C52te] Current Opinion in Obstetrics and Gynecology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Onco

[C53te] Current Opinion in Oncology, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Opht

[C54te] Current Opinion in Ophthalmology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Orth

[C55te] Current Opinion in Orthopaedics, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Pedi

[C56te] Current Opinion in Pediatrics, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Psyc

[C57te] Current Opinion in Psychiatry, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Radi

[C58te] Current Opinion in Radiology, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Rheu

[C59te] Current Opinion in Rheumatology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Opin-Stru-Biol

[C60te] Current Opinion in Structural Biology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Pate-Fast-Onli-Acce

[C61te] Current Patents Fast-Alert— Online Access, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Prac-Medi

[C62te] Current Practice of Medicine, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Curr-Ther-Pate

[C63te] Current Therapeutic Patents, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Dete-Mana-Bloo-Lipi

[C64te] Detection and Management of Blood Lipid Disorders, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Euro-Jour-Gast-Hepa

[C65te] European Journal of Gastroenterology and Hepatology, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Gast-Elde

[C66te] Gastroenterology in the Elderly, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Hand-Hear-Fail

[C67te] Handbook of Heart Failure, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Hand-High-Bloo-Pres

[C68te] Handbook of High Blood Pressure, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Hand-Hype

[C69te] Handbook of Hyperlipidaemia, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Herb-Fung-Plan-Grow-Regu

[C70te] Herbicide, Fungicide and Plant Growth Regulator Patent Fast-Alert, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Hype-Diab

[C71te] Hypertension and Diabetes, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Hype-Elde

[C72te] Hypertension and the Elderly, 1987–date. Current Science, 34– 42 Cleveland Street, London, UK. See [C01te].

CurSci:Hype-Annu-1989–90

[C7389] Hypertension Annual 1989–90, 1989. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Hype-Annu-1990–91

[C7490] Hypertension Annual 1990–91, 1990. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Immu

[C75te] Immunobiology, 1987–date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Jour-Hype

[C76te] Journal of Hypertension, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Lect-Hype

[C77te] Lectures in Hypertension, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Macr-Stru-1990

[C7890] Macromolecular Structures 1990, 1990. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Pest-Inse-Fert-Pate-Fast

[C79te] Pesticide, Insecticide and Fertilizer Patent Fast-Alert, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Pock-Guid-Acne

[C80te] Pocket Guide to Acne, 1987– date. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Slid-Atla-Gast-Endo-1990

[C8190] Slide Atlas of Gastrointestinal Endoscopy 1990, 1990. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Slid-Atla-Gast-Endo-1991

[C8291] Slide Atlas of Gastrointestinal Endoscopy 1991, 1991. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Slid-Atla-Curr-Opht-1990

[C8390] Slide Atlas of Current Ophthalmology 1990, 1990. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Slid-Atla-Curr-Opht-1991

[C8491] Slide Atlas of Current Ophthalmology 1991, 1991. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Slid-Atla-Curr-Radi-1990

[C8590] Slide Atlas of Current Radiology 1990, 1990. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

CurSci:Slid-Atla-Curr-Radi-1991

[C8691] Slide Atlas of Current Radiology 1991, 1991. Current Science, 34–42 Cleveland Street, London, UK. See [C01te].

Calingaert:PTF88

[Cal88]  Peter Calingaert.  Program

Translation Fundamentals—Methods and Issues. Computer Science Press, 1803 Research Boulevard, Rockville, MD 20850, USA, 1988. ISBN 0-88175-096-4. xvii + 366 pp. LCCN QA76.76.T83 C35 1988.

Carberry:PRN90

[Car90] Sandra Carberry. Plan Recognition in Natural Language Dialogue. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-03167-1. xi + 286 pp. LCCN Q335 .C28 1990.

Cashdan:RUT90

[Cas90] Elizabeth A. Cashdan, editor.

Risk and Uncertainty in Tribal and Peasant Economies. Westview Press, 5500 Central Ave., Boulder, CO 80301-2847, USA, 1990. ISBN 0-8133-7821-4. x + 342 pp. LCCN GN448.2 .R57 1990.

Cheswick:2003:FIS

[CBR03] William R. Cheswick, Steven M. Bellovin, and Aviel D. Rubin. Firewalls and Internet Security: Repelling the Wily Hacker. Addison-Wesley, Reading, MA, USA, second edition, 2003. ISBN 0-201-63466-X. xiv + 433 pp. LCCN TK5105.875.I57C44 2003. US$49.99, CAN$77.99. Prepared with L^{A}TEX.

Curry:DSF89

[CDF89] Guy L. Curry, Bryan L. Deuermeyer, and Richard M. Feldman. Discrete Simulation: Fundamentals and Microcomputer Support. Holden-Day, Inc., Oakland, CA 94609, USA, 1989. ISBN 0-8162-2060-3.

Carriero:HWP90

[CG90] Nicholas Carriero and David Gelernter. How to Write Parallel Programs. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-03171-X. 232 pp. LCCN QA76.642 .C37 1990.

Char:1991:MLVb

[CGG^{+}91a] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. Maple Library V Reference Manual. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97592-6, 3-540-97592-6. xxv + 698 pp. LCCN QA155.7.E4 M353 1991.

Char:1991:MVLa

[CGG^{+}91b] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. Maple V: Language Reference Manual. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97622-1 (New York), 3-540-97622-1 (Berlin). xv + 267 pp. LCCN QA155.7.E4 M36 1991. US$24.95, FF 193,00.

Char:1992:FLT

[CGG^{+}92] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. First Leaves: a Tutorial Introduction to Maple V. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97621-3, 3-540-97621-3. xvii + 253 pp. LCCN QA155.7.E4 F56 1992. Also available in Japanese, ISBN 4-431-70651-8.

Chambers:SMS93

[CH93] John M. Chambers and Trevor J.

Hastie, editors. Statistical Models in S. Chapman and Hall, Ltd. and Wadsworth and Brooks/Cole Publishing Company, London, UK and Pacific Grove, CA, USA, 1993.

ISBN 0-412-05301-2 (Chapman and Hall), 0-534-16765-9 (Wadsworth paperback), 0-534-16764-0 (Wadsworth hardcover). xv + 608 pp. LCCN QA276.4 .S65 1992. US$40.95.

Chapman:VIA91

[Cha91] David Chapman. Vision, Instruction, and Action. The MIT Press, Cambridge, MA, 1991. ISBN 0-262-03181-7. ix + 295 pp. LCCN Q335 .C4829 1991. Prepared with L^{A}TEX.

Clark:CRU84

[Cla84] Malcolm Clark, editor. Coastal research: UK perspectives. Geo Books, Norwich, UK, 1984. ISBN 0-86094-166-3. LCCN GB457.21 .C61 1984. Typeset using TEX80.

Clark:PPF86

[Cla86] Malcolm Clark. pc-Portable Fortran. Computers and their applications. Ellis Horwood, New York, NY, USA, 1986. ISBN 0-7458-0005-X. 228 pp. LCCN QA76.73.F25 C55 1986. Laserprinted using TEX.

Cormen:IA89

[CLR89] Thomas Cormen, Charles Leiserson, and Ronald Rivest. Introduction to Algorithms. The MIT Press, Cambridge, MA, 1989. ISBN 0-262-03141-8. xvii + 1028 pp. LCCN QA76.6 .C662 1989.

Chandy:PPD88

[CM88] K. Mani Chandy and Jayadev Misra. Parallel Program Design: A Foundation. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-05866-9. xxviii + 516 pp. LCCN QA76.6 .C42818 1988b. Prepared with L^{A}TEX.

CMUC

[CMU89] Commentationes Mathematicae Universitatis Carolinae, 1989. Publication of the Faculty of Mathematics and Physics, Charles University, Prague, Czechoslovakia. This was prepared with AMS-TEX.

Conrad:FPM89

[Con89] Lawrence I. Conrad. The Formation and Perception of the Modern Arab World: Studies by Marwan R. Buheiry. The Darwin Press, Princeton, NJ, USA, 1989. ISBN 0-87850-064-2. 624 pp. LCCN DS63.2.U5 B82 1989. Prepared with L^{A}TEX. PCTEX and Computer Modern fonts, with final output on an HP LaserJet II at 300dpi.

cornishman

[Corte] Peter Abbott, editor. The Cornishman, 1989–date. ISSN 0964-5098. Gloucestershire Warwickshire Railway Ltd., The Railway Station, Toddington, Gloucestershire, GL54 5DT, UK. Prepared with L^{A}TEX.; incorporates numerous photographs and illustrations as EPS files. From issue 42 (Autumn 1992), photographs will be stored in compressed TIFF format and converted to PostScript ‘on the fly’.

Craig:IRM89

[Cra89] John J. Craig. Introduction to Robotics: Mechanics and Control. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-09528-9. xiii + 450 pp. LCCN TJ211.C67 1989. Prepared with TEX.

Coyne:KBD90

[CRR^{+}90] R. D. Coyne, M. A. Rosenman, A. D. Radford, M. Balachandran, and J. S. Gero. Knowledge-Based Design Systems. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-10381-8. xi + 567 pp. LCCN TA174.K59 1990. Prepared with TEX.

Complex-Systems

[CS87] Complex Systems. Complex Systems Publications Inc., 1987. ISSN 0891-2513. Technical journal published six times per year, using Computer Modern fonts.

Cuoco:C89

[Cuo89] Albert Cuoco. Investigations in Algebra. The MIT Press, Cambridge, MA, 1989. ISBN 0-262-53071-6 (paper), 0-262-03144-2 (hard). 623 pp. LCCN QA159 .C86 1990.

Cavalotto:SMM90

[CV90] M. Cavalotto and F. Vinardi. I sistemi meccanici, Modelli e analisi funzionale. Edizioni Cooperativa Libraria Universitaria Torino, Corso Duca degli Abruzzi, 24, I 10129 Torino, Italy, 1990. Prepared with L^{A}TEX.

Chung:ISI83

[CW83] K. L. Chung and R. J. Williams.

Introduction to Stochastic Integration. Birkhäuser, 1983. ISBN 0-8176-3117-8, 3-7643-3117-8. xiii + 191 pp. LCCN QA274.22 .C48 1983.

Dai:UPT90

[Dai90]  Catherine Dai.  Under the

Phoenix Tree. Bookman, Taipei, Taiwan, 1990. ISBN 957-586-063-2. 330 pp. LCCN PL2855.A3 U54 1990.

Daigle:QTC91

[Dai91] John W. Daigle. Queuing Theory for Computer Communications. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-06755-2. xvi + 414 pp. LCCN TK5105.5.D325 1991. Prepared with TEX.

Dai:SCT92

[Dai92] Catherine Dai. Bound Feet: Stories of Contemporary Taiwan. Bookman, Taipei, Taiwan, 1992. ISBN 957-586-241-4.

Dally:VLS87

[Dal87] W. J. Dally. A VLSI Architecture for Concurrent Data Structures. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1987. ISBN 0-89838-235-1. xvi + 243 pp. LCCN TK7888.4 .D341 1987. Prepared with TEX (unknown variety).

deBroglie:HUP90

[dB90]  Louis de Broglie.  Heisen-

berg’s Uncertainty Principle and the Probabilistic Interpretation of Wave Mechanics. Number 40 in Fundamental Theories of Physics. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0929-4. xliv + 302 pp. LCCN QC174.2 .B653513 1990. English translation by Alwyn van der Merwe. Prepared with TEX.

Dongarra:1991:SLS

[DDSv91] Jack J. Dongarra, Iain S. Duff, Danny C. Sorensen, and Henk A. van der Vorst. Solving Linear Systems on Vector and Shared Memory Computers. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991. ISBN 0-89871-270-X. x + 256 pp. LCCN QA184 .S65 1991.

DamEngineering

[DE90] Dam engineering, January 1990. ISSN 0958-9341. Laser-printed using L^{A}TEX.

Devroye:1986:LNB

[Dev86] Luc Devroye. Lecture Notes on Bucket Algorithms. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1986. ISBN 0-8176-3328-6. 148 pp. LCCN QA76.9.D35 D48 1986.

Devaney:ICD87

[Dev87] Robert L. Devaney. An Introduction to Chaotic Dynamical Systems. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-8053-1601-9. xii + 320 pp. LCCN QA614.8 .D48 1986. Prepared with TEX.

Devaney:ICD89

[Dev89] Robert L. Devaney. An Introduction to Chaotic Dynamical Systems. Addison-Wesley, Reading, MA, USA, second edition, 1989. ISBN 0-201-13046-7. xvi + 336 pp. LCCN QA614.8 .D48 1989. Prepared with TEX.

Devito:FAL90

[Dev90] Carl L. Devito. Functional Analysis and Linear Operator Theory. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-11941-2. x + 358 pp. LCCN QA329.2 .D48 1990. Prepared with L^{A}TEX.

Doolen:LGM90

[DFH^{+}90] Gary D. Doolen, M. Frisch, B. Hasslacher, S. Oszag, and S. Wolfram, editors. Lattice Gas Methods for Partial Differential Equations. Addison-Wes-ley, Reading, MA, USA, 1990. ISBN 0-201-16579-1 (hardcover), 0-201-13232-X (paperback). xix + 554 pp. LCCN QA374 .L34 1989. Prepared with TEX.

Dibble:OIA92

[Dib92] Peter C. Dibble. OS-9 Insights —An Advanced Programmers Guide to OS-9. Microware Systems Corporation, Des Moines, IA, USA, second edition, 1992. ISBN 0-918035-03-1. Prepared with L^{A}TEX. Used a private modification of the L^{A}TEX book format with Adobe Garamond PostScript type faces. dvips produced 300dpi PostScript output that was printed on a Hewlett-Packarad LaserJet II with Adobe’s PostScript cartridge. Final copy was produced from PostScript disks sent to Courier Connection who returned pallets of books in something like four weeks. The titlepages are hand-coded Postscript, and figures use the figure environment.

Dijkstra:FDP89

[Dij89] Edsger W. Dijkstra, editor. Formal Development of Programs and Proofs. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-17237-2. xiv + 242 pp. LCCN QA76.6 .F657 1990. Prepared with L^{A}TEX.

Davis:KBS82

[DL82] R. Davis and D. B. Lenat. Knowledge-Based Systems in Artificial Intelligence. McGraw-Hill, New York, NY, USA, 1982. ISBN 0-07-015557-7. xxi + 490 pp. LCCN Q335 .D38. Prepared with TEX.

Doughty:LII90

[Dou90] N. A. Doughty. Lagrangian interaction: an introduction to relativistic symmetry in electrodynamics and gravitation. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-41625-5. xix + 569 pp. LCCN QC631.D68 1989. Prepared with L^{A}TEX.

Desaubies:OGT90

[DTZJ90] Yves Desaubies, Albert Tarantola, and Jean Zinn-Justin, editors. Oceanographic and Geophysical Tomography. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1990. Les Houches Session L (1989).

Demana:CAG89

[DW89a] Franklin Demana and Bert K.

Waits. College Algebra: A Graphing Approach. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-52811-8. LCCN QA155.5 .D45 1989. Prepared with TEX.

Demana:CAT89

[DW89b] Franklin Demana and Bert K.

Waits. College Algebra and Trigonometry: A Graphing Approach. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-52810-X. xvi + 637 + 67 + 7 pp. LCCN QA331.3 .D46 1991. Prepared with TEX.

Demana:TGA89

[DW89c] Franklin Demana and Bert K.

Waits. Trigonometry: A Graphing Approach. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-52812-6. xii + 482 pp. LCCN QA538 .D45 1990. Prepared with TEX.

Demana:PFG90

[DW90a] Franklin Demana and Bert K.

Waits. Precalculus: Functions and Graphs. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-52781-2. 718 + 95 + 16 pp. LCCN QA154.2 .D43 1990. Prepared with TEX.

Demana:PM90

[DW90b] Franklin Demana and Bert K. Waits. Precalculus Mathematics: A graphing approach. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-52813-4. xvi + 933 pp. LCCN QA109 .D45 1990. Prepared with TEX.

Dybvig:SPL87

[Dyb87] R. Kent Dybvig. The SCHEME Programming Language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-791864-X (paperback). xi + 242 pp. LCCN QA76.73.S34 D93 1987.

Epstein:RTC90

[ES90] Lawrence Epstein and Richard F. Sherburne. Reflections on Tibetan Culture: Essays in Memory of Turrell V. Wylie, volume 12 of Studies in Asian Thought and Religion. Mellen Press, Lewiston, NY, USA, 1990. ISBN 0-88946-064-7. xiii + 350 pp. LCCN DS786 .R43 1990.

Estrade:SJR90

[Est90] J. Estrada. Los Sonidos en Juan Rulfo. Universidad Autonoma de México, México DF, México, 1990.

Ellis:TTP90-1

[ET90] Paul J. Ellis and Y. C. Tang, editors. Trends in Theoretical Physics, volume 1. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-50393-X. xv + 412 pp. LCCN QC20 .T69 1990 Sci-Eng. Prepared with TEX.

Ellis:TTP91-2

[ET91] Paul J. Ellis and Y. C. Tang, editors. Trends in Theoretical Physics, volume 2. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-52251-9. LCCN QC20 .T69 1990. Prepared with TEX.

tj-elek-vestnik

[EVxx] Elektrotehniški Vestnik, 19xx.

Typeset with LaTeX on a Apple Macintosh II, with final pages produced on a 300dpi Apple LaserWriter printer, soon to be replaced by a PostScript phototypesetter. Illustrations done by paste-up, but will soon be handled electronically. Articles are published in Slovenian, Serbo-Croatian, and English in Electrical Engineering and Computer Science. Authors are encouraged to submit manuscripts using the standard L^{A}TEX article.sty file.

Feferman:KG86

[FDK^{+}86] Solomon Feferman, John W.

Dawson, Stephen C. Kleene, Gregory H. Moore, Robert M. Solovay, and Jean van Heijenoort. Kurt G̈odel: Collected Works. Vol. 1: Publications 1929-1936. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1986. ISBN 0-19-503964-5. 459 pp. LCCN QA9 .G53131 1986. US$35. From Computing Reviews: “The book is very well written and put together as if by a magic amalgamation of C. P. Snow and A. Church. Following the publication style on Wittgenstein’s works, there are accurate, facingpage English translations of the original German texts. All type has been reset by computer in a more readable form. Indeed, the sheer joy of TEX blooms forth on every page. Every school library should contain this definitive series of volumes.”.

Friedman:LL87

[FF87] Daniel Friedman and Matthias Felleisen. The Little LISPer. The MIT Press, Cambridge, MA, 1987. ISBN 0-262-56038-0. xiv + 186 pp. LCCN QA76.73.L23 F74 1987.

Fraser:1995:RCC

[FH95] Chris W. Fraser and David R. Hanson. A Retargetable C Compiler: Design and Implementation. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, 1995. ISBN 0-8053-1670-1. xv + 564 pp. LCCN QA76.73.C15F75 1995. US$55.95. URL https:/

/www.pearson.com/us/higher-
education/program/Hanson-
Retargetable-C-Compiler-A-
Design-and-Implementation/
PGM166351.html. From the au-

thors’ announcement: “… this book is an example of a ‘literate program.’ Like TEX: The Program by D. E. Knuth (Addison-Wesley, 1986), this book is lcc’s source code and the prose that describes it. The code is presented in the order that best suits understanding, not in the order dictated by the C programming language. The source code that appears on the diskette is extracted automatically from the book’s text files. … The distribution is available via ‘anonymous’ ftp from

ftp.cs.princeton.edu  (128.

112.152.13) in the directory pub/lcc. … Additional information about lcc and about our book is available on the

WWW at URL http://www.
cs.princeton.edu/software/
lcc.”.

Field:APQ89

[Fie89] Richard D. Field. Applications of Perturbative QCD. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-16643-7. Prepared with TEX.

Flynn:1995:WHH

[Fly95]  Peter Flynn.  The World-

WideWeb Handbook: An HTML Guide for Users, Authors and Publishers. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1995. ISBN 1-85032-205-8. xix + 351 pp. LCCN TK5105.888 .F56 1995. Foreword by Tim Berners-Lee. Three sections deal with (1) Getting connected to the Internet and using Internet software; (2) Writing HTML (2.0) files for the World-WideWeb; (3) Running a HTTP server and providing a Web service. Author is a member of the IETF Working Group on HTML.

Text includes additional material on SGML; choice of editors, browsers and servers; copyright and intellectual property; and advance details of HTML3.

Fortier:HLT89

[For89] Paul J. Fortier. Handbook of LAN Technology. McGraw-Hill, New York, NY, USA, 1989. ISBN 0-07-021623-1. xxiv + 551 pp. LCCN TK5105.7.F67 1989. Prepared with TEX.

Franklin:FCD91

[FPEN91] Gene F. Franklin, J. David Powell, and Abbas Emami-Naeini. Feedback Control of Dynamic Systems. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-50862-1. xvi + 672 pp. LCCN TJ216 .F723 1991 Reserve. Prepared with TEX.

Franklin:DCD90

[FPW90] Gene F. Franklin, David J. Powell, and Michael L. Workman. Digital Control of Dynamic Systems. Addison-Wesley, Reading, MA, USA, second edition, 1990. ISBN 0-201-11938-2. xxi + 837 pp. LCCN TJ223.M53 F73 1990. Prepared with AMS-TEX.

Falck:TM89

[FR89] Willie Falck and Johan Hammond Rosbach. Tra la mondo (All around the world). Esperantoforlaget AS, Oslo, Norway, 1989. ISBN 82-90006-07-1. Prepared with L^{A}TEX. This is a guidebook for teaching Esperanto.

Fradkin:HTS91

[Fra91] Eduardo Fradkin. High Temperature Superconductivity and Strongly Correlated Electron Systems. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-52253-5. xvi + 350 pp. LCCN QC611.98.H54 F73 1991. Prepared with TEX.

Freeman:NN91

[FS91] James A. Freeman and David M.

Skapura. Neural Networks. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-51376-5. xiii + 401 pp. LCCN QA76.87.F74 1991. Prepared with L^{A}TEX.

Feijen:BOB90

[FvGGM90] W. H. J. Feijen, A. J. M. van Gasteren, D. Gries, and J. Misra, editors. Beauty is our business: a birthday salute to Edsger W. Dijkstra. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-97299-4. xix + 453 pp. LCCN QA76 .B326 1990. Prepared with L^{A}TEX.

Fox:PCW94

[FWM94] Geoffrey C. Fox, Roy D.

Williams, and Paul C. Messina. Parallel Computing Works! Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1994. ISBN 1-55860-253-4. xviii + 977 pp. LCCN QA76.58.F68 1994. US$64.95. Prepared with L^{A}TEX.

Goodheart:MGE94

[GC94] Berny Goodheart and James Cox. The Magic Garden Explained: The Internals of UNIX System V Release 4, an Open Systems Design. Prentice-Hall, 1994. ISBN 0-13-098138-9. xxviii + 664 pp. LCCN QA76.76.O63 G6633 1994. US$38.00. Probably a good companion to Bach’s book, The Design of the UNIX Operating System, …. Covering the internals, system calls, kernal of System V Release 4 ….

Gazdar:NLP87

[GFOE87] Gerald Gazdar, Alex Franz, Karen Osborne, and Roger Evans. Natural Language Processing in the 1980s, A Bibliography, volume 12 of Center for the Study of Language and Information Notes. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1987. ISBN 0-937073-26-1 (hardcover), 0-937073-28-8 (paperback). viii + 240 pp. LCCN P98.A12 N271 1987. US$24.95 (hardcover), US$9.95 (paperback). This book was prepared automatically from a database and set with TEX. The intent is to keep it updated.

Gander:SPS93

[GH93] Walter Gander and Jiři Hřebíček. Solving Problems in Scientific Computing Using MAPLE and MATLAB. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-57329-1, 3-540-57329-1. xiii + 268 pp. LCCN Q183.9 .G36 1993. Prepared with L^{A}TEX.

Gunion:HHG90

[GHKD90] John F. Gunion, Howard E.

Haber, Gordon L. Kane, and Sally Dawson. The Higgs Hunter’s Guide. Addison-Wes-ley, Reading, MA, USA, 1990. ISBN 0-201-50935-0. xvi + 425 pp. LCCN QC793.5.B62 H54 1990. Prepared with TEX.

Gibson:PNA90

[Gib90] Rosalind S. Gibson. Principles of Nutritional Assessment. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-505838-0. xvi + 691 pp. LCCN RC621 .G52 1990. Set by L^{A}TEX, with graphics from Harvard Graphics and Adobe Illustrator, and designed by Ian L. Gibson and Philip Taylor with the assistance of the publisher. The book is typeset in Adobe Times Roman 10.5/12 and 8.5/10.5, with figure annotation in Adobe Helvetica. The reasons for the unusual font size are described in Philip Taylor’s TEX91 conference paper, to appear in TUGboat 12, no. 1, 1991.

Girard:RDC90

[Gir90] André Girard. Routing and Dimensioning in Circuit-Switched Networks. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-12792-X. xiv + 556 pp. LCCN TK5103.8.G57 1990. Prepared with INRSTEX.

Greene:MAA81

[GK81] Daniel H. Greene and Donald E. Knuth. Mathematics For the Analysis of Algorithms. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1981. ISBN 3-7643-3046-5. 107 pp. LCCN QA76.6 .G7423 1981. US$10.00.

Greene:1982:MAA

[GK82] Daniel H. Greene and Donald E. Knuth. Mathematics for the Analysis of Algorithms. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, second edition, 1982. ISBN 3-7643-3102-X. 123 pp. LCCN QA76.6 .G7423 1982.

Gavrilov:EAI92

[GKL^{+}92] A. K. Gavrilov, I. V. Kuklina, I. A. Levinskaja, A. A. Nejjkhardt, and I. A. Shishova. Ehtudy po antichnojj istorii i kulqture Severnogo Prichernomorqja (English: Studies on History and Culture of the North Coast of the Black Sea in Antique Time). Glagol, St. Petersburg, Russia, 1992. ISBN 5-85381-002-2. 271 pp.

Graham:1989:CM

[GKP89] Ronald L. Graham, Donald E. Knuth, and Oren Patashnik. Concrete Mathematics. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-14236-8. xiii + 625 pp. LCCN QA39.2 .G7331 1989.

Graham:1994:CM

[GKP94] Ronald L. Graham, Donald E. Knuth, and Oren Patashnik.

Concrete Mathematics. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-55802-5. xiii + 657 pp. LCCN QA39.2 .G733 1994.

Gill:PO81

[GMW81] Philip E. Gill, Walter Murray, and Margaret H. Wright. Practical Optimization. Academic Press, Boston, MA, USA, 1981. ISBN 0-12-283950-1 (hardcover), 0-12-283952-8 (paperback). xvi + 401 pp. LCCN QA402.5 .G5 1981.

Gill:NLA91-1

[GMW91] P. Gill, E. Murray, and M. Wright.

Numerical Linear Algebra and Optimization, volume 1. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-12649-4. LCCN QA 184 G55 1991. Prepared with TEX.

Golub:1992:SCD

[GO92] Gene H. Golub and James M. Ortega. Scientific Computing and Differential Equations: An Introduction to Numerical Methods. Academic Press, New York, NY, USA, second edition, 1992. ISBN 0-12-289255-0. xi + 337 pp. LCCN QA371 .G62 1992. Prepared with L^{A}TEX.

Grimson:AIB87

[GP87] W. E. L. Grimson and R. S.

Patil, editors. A. I. in the 1980s and Beyond: An MIT Survey. The MIT Press, Cambridge, MA, 1987. ISBN 0-262-07106-1. 374 pp. LCCN Q335 .A42 1987.

Prepared with Daniel Brotsky’s YTEX.

Granshaw:HAO88

[Gra88] Lindsay Granshaw. Health for All’: The origins of the National Health Service, 1848–1948. A Fortieth Anniversary Retrospect. Wellcome Institute for the History of Medicine, London, UK, 1988. ISBN 0-85484-059-1. Prepared with L^{A}TEX. PCTEX and Times Roman fonts, with James Clark’s dvitops for final output on bromide at 1270dpi on the ULCC Linotronic 300.

Grandine:1990:NMP

[Gra90a] Thomas A. Grandine. The Numerical Methods Programming Projects Book. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853385-3 (hardcover), 0-19-853387-X (paperback). viii + 146 pp. LCCN QA76.6 .G718 1990.

Gray:T90

[Gra90b] Alfred Gray. Tubes. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-15676-8. xii + 283 pp. LCCN QC174.27.W28 G73 1990. Prepared with L^{A}TEX.

Greene:HGA91

[Gre91] Andrew M. Greene, Ed. How to Get Around MIT. Technology Community Association, Cambridge, MA, USA, twenty-third edition, 1991. Press run of 3000 copies; sold in local bookstore, and distributed gratis to 2000 students. Typeset with Plain TEX, using Computer Modern fonts, with markup macros written in-house. perl was used to provide a front end, consistent pagination, and the index. Final pages were prepared on 300dpi Apple LaserWriter Plus printer.

Grimson:FIS81

[Gri81] W. Eric L. Grimson. From Images to Surfaces: A Computational Study of the Human Early Visual System. The MIT Press, Cambridge, MA, 1981. ISBN 0-262-07083-9. 274 pp. LCCN QP487 .G74. Prepared with TEX.

Gries:EMR89

[Gri89] David Gries, editor. Encapsulation, Modularization, and Reusability. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-17233-X. Prepared with L^{A}TEX.

Grimson:ORC91

[Gri91] W. Eric L. Grimson. Object Recognition by Computer: The Role of Geometric Constraints. The MIT Press, Cambridge, MA, 1991. ISBN 0-262-07130-4. xv + 512 pp. LCCN TA1632 .G735 1990. Prepared with Daniel Brotsky’s YTEX.

Gaarding:ACS88

[GT88a] L. Gårding and T. Tambour.

Algebra for Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96780-X. ix + 198 pp. LCCN QA76.9.M35 G37 1988.

Gould:ICS88

[GT88b] Harvey Gould and Jan Tobochnik. Introduction to Computer Simulation Methods: Applications to Physical Systems, Part 1 and 2. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-16503-1 (hardcover), 0-201-16504-X (paperback). LCCN QC21.2 .G67 1988. Prepared with TEX.

Gudder:QP88

[Gud88] Stanley P. Gudder. Quantum probability. Probability and mathematical statistics. Academic Press, Boston, MA, USA, 1988. ISBN 0-12-305340-4. xii + 316 pp. LCCN QC20.7.P7 G83 1988. Prepared with TEX.

Gudder:MJ94

[Gud94] Stanley P. Gudder. A Mathematical Journey. McGraw-Hill, New York, NY, USA, second edition, 1994. ISBN 0-07-025130-4. LCCN QA39.2 .G83 1994. US$42.50. Prepared with TEX.

Gutowitz:CAT90

[Gut90] Howard Gutowitz, editor. Cellular Automata: Theory and Experiment. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1990. ISBN 0-262-57086-6. LCCN QA267.5.C45 C46 1991. Published as Physica D45 (1990) Nos. 1-3.

Golub:1989:MC

[GV89] Gene H. Golub and Charles F. Van Loan. Matrix Computations, volume 3 of Johns Hopkins Series in the Mathematical Sciences. The Johns Hopkins University Press, Baltimore, MD, USA, second edition, 1989. ISBN 0-8018-3772-3 (hardcover), 0-8018-3739-1 (paperback). xix + 642 pp. LCCN QA188 .G65 1989. US$14.50.

Halley:THT88

[Hal88] J. Woods Halley, editor. Theories of High Temperature Superconductivity. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-12008-9. vi + 264 pp. LCCN QC612.S8 T437 1988. Prepared with TEX.

Hammouche:AVA90

[Ham90] Rabah Hammouche. Atlas Vent de l’Algérie. Office National de la Météorologie, Alger, Algérie, 1990. 150 pp. Typeset with Bitstream fonts on a Canon Series III 300-dpi laser printer, with extensive graphics and data tables; the data files are available on IBM PC diskettes.

Hamming:APS91

[Ham91] Richard W. Hamming. The Art of Probability—For Scientists and Engineers. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-51058-8. xvii + 344 pp. LCCN QA273 .H3544 1991.

Harvey:CSL85

[Har85] Brian Harvey. Computer Science Logo Style, volume 1. The MIT Press, Cambridge, MA, 1985.

ISBN 0-262-58072-1. LCCN QA76.6.H385 1985.

Harvey:CSL86

[Har86] Brian Harvey. Computer Science Logo Style, volume 2. The MIT Press, Cambridge, MA, 1986. ISBN 0-262-58080-2. LCCN QA76.6.H385 1985.

Harvey:CSL87

[Har87] Brian Harvey. Computer Science Logo Style, volume 3. The MIT Press, Cambridge, MA, 1987. ISBN 0-262-58086-1.

Hatfield:QFT91

[Hat91] Brian F. Hatfield. Quantum Field Theory of Point Particles and Strings. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-11982-X. xvii + 734 pp. LCCN QC 174.45 H29 1992. Prepared with TEX.

Horn:SS89

[HB89] Berthold K. P. Horn and Michael J. Brooks. Shape from Shading. The MIT Press, Cambridge, MA, 1989. ISBN 0-262-08183-0. 577 pp. LCCN TA1632 .S53 1989. Prepared with Daniel Brotsky’s YTEX.

Heck:IM93

[Hec93] André Heck. Introduction to Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97662-0, 3-540-97662-0. xiii + 497 pp. LCCN QA155.7.E4H43 1993.

Helfer:IMD90

[Hel90] Adam Helfer. Introduction to Modern Differential Geometry. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-50061-2. Prepared with TEX.

Huttner:RAE93

[HH93] Sidney F. Huttner and Elizabeth Stege Huttner. A register of artists, engravers, booksellers, bookbinders, printers and publishers in New York City, 1821– 42. The Bibliographical Society of America, New York, NY, USA, 1993. ISBN 0-914930-15-X. 299 pp. LCCN Z478.6.N5 H87 1993. US$50.00. Prepared with TEX.

Higham:1996:ASN

[Hig96] Nicholas J. Higham. Accuracy and Stability of Numerical Algorithms. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-355-2 (paperback). xxviii + 688 pp. LCCN QA297.H53 1996. US$39.00. URL http:

//www.ma.man.ac.uk/~higham/
asna.html.

Hildreth:MVM84

[Hil84] E. C. Hildreth. The Measurement of Visual Motion. ACM Distinguished Dissertation Series. The MIT Press, Cambridge, MA, 1984. ISBN 0-262-08143-1. 241 pp. LCCN BF241 .H53 1984. Prepared with Daniel Brotsky’s YTEX.

Hillis:CM85

[Hil85] W. Daniel Hillis. The Connection Machine. The MIT Press, Cambridge, MA, 1985. ISBN 0-262-08157-1. xiii + 190 pp. LCCN QA267 .H4871 1985. The book was set using L^{A}TEX by Thinking Machines Corporation and printed and bound by Halliday Lithograph in the USA. The terms in the index were selected automatically using The TMC Indexer.

Hill:GPL94

[Hil94] Patricia Hill. The Gödel Programming Language. The MIT Press, Cambridge, MA, 1994. ISBN 0-262-08229-2. xx + 348 pp. LCCN QA76.73.G17H55 1994.

Hiscocks:1996:DLA

[His96] Richard D. Hiscocks. Design of Light Aircraft. Richard D. Hiscocks, Apt 704, Beach Park Apartments, 2095 Beach Avenue, Vancouver, British Columbia V6G 1Z3, Canada; Tel: (604) 682-6214, 1996. ISBN 0-9699809-0-6. 281 pp. LCCN TL671.2.H57 1995. US$29.89, CDN$39.68.

Hansen:LFG91

[HLZJ91] J. P. Hansen, D. Levesque, and J. Zinn-Justin, editors. Liquids, Freezing and Glass Transition (parts I, II). Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1991. Les Houches Session LI (1990).

Hecht-Nielsen:N90

[HN90] Robert Hecht-Nielsen. Neurocomputing. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-09355-3. xiii + 433 pp. LCCN QA76.5.H4442 1990. Prepared with L^{A}TEX.

Hoare:DCC91

[Hoa91] C. A. R. Hoare. Developments in Concurrency and Communication. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-17232-1. xiv + 336 pp. LCCN QA76.58 .D48 1991. Prepared with L^{A}TEX.

Hoenig:AFM90

[Hoe90] Alan Hoenig. Applied Finite Mathematics. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-557093-9. xiv + 619 pp. LCCN QA39.2 .H6227 1989. US$25.00.

Holmang:IUS88

[Hol88]  Per Olof Holmäng.  Idrott

och Utrikespolitk: Den svenska idrottsrörelsens internationella förbindelser 1919–1945. Number 37 in Meddelande från Historiska Institutionen i Göteborg. University of Gothenburg Press, Gothenburg, Sweden, 1988. ISBN 91-7900-635-3. ISSN 0346-914X. 329 pp. LCCN GV706.35 .H65 1988. Title in English: The international relations of the Swedish sports movement 1919– 1945 (Publications of the Department of History, University of Gothenburg), no. 37.

Horn:RV86

[Hor86] Berthold K. P. Horn. Robot Vision. The MIT Press, Cambridge, MA, 1986. ISBN 0-262-08159-8. x + 509 pp. LCCN TJ211.3 .H671 1986. Prepared with Daniel Brotsky’s YTEX.

Hertz:NC91-1

[HPK91] J. Hertz, R. Palmer, and A. Krogh. Neural Computation, volume 1. Addison-Wes-ley, Reading, MA, USA, 1991. ISBN 0-201-50395-6 (hardcover), 0-201-51560-1 (paperback). xxii + 327 pp. LCCN QA76.5 .H475 1991. Prepared with TEX.

Herrmann:SMF90

[HR90] Hans J. Herrmann and Stéphan Roux. Statistical Models for the Fracture of Disordered Media. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1990. ISBN 0-444-88551-X (hardcover), 0-444-88550-1 (paperback). xiii + 353 pp. LCCN TA409 .S73 1990.

Haralick:CRV91-1

[HS91] Robert M. Haralick and Linda D. Shapiro. Computer and Robot Vision, volume 1. Addison-Wes-ley, Reading, MA, USA, 1991. ISBN 0-201-10877-1. LCCN TA1632 .H37 1992. Prepared with TEX.

Huet:LFF89

[Hue89] Gérard Huet, editor. Logical Foundations of Functional Programming. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-17234-8. xvi + 491 pp. LCCN QA76.62 .L64 1990. Prepared with L^{A}TEX.

Husek:RPG92

[HvM92] Miroslav Hušek and Jan van Mill, editors. Recent Progress in General Topology. North-Holland Publishing Company, Amsterdam, The Netherlands, 1992. ISBN 0-444-89674-0. xi + 796 pp. LCCN QA611.A1 R43 1992. Prepared with L^{A}TEX. K. P. Hart developed a L^{A}TEX style for this book modelled on North-Holland conventions.

Hairer:SOD91

[HW91] Ernst Hairer and Gerhard Wanner. Solving ordinary differential equations II: Stiff and differential-algebraic problems, volume 14 of Springer series in computational mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-53775-9. LCCN QA372 .H16 1987. Prepared with TEX.

Henzler:OF91

[HwCZ91] Martin Henzler and Wolfgang Göpel with Christiane Ziegler. Oberflächenphysik des Festkörpers. B. G. Teubner, Stuttgart, Germany, 1991. ISBN 3-519-03047-0. Prepared with L^{A}TEX.

Harper:GCA91

[HWH91] David Harper, Chris Wooff, and David Hodgkinson. A Guide to Computer Algebra Systems. John Wiley, New York, NY, USA, 1991. ISBN 0-471-92910-7. xii + 148 pp. LCCN QA155.7.E4 H37 1991. Prepared with L^{A}TEX.

IEEE:el-ins

[IEE89] Ieee transactions on electrical insulation, 1989. ISSN 0018-9367.

Jain:OSI90

[JA90] Bijendra N. Jain and Ashok K. Agrawala. Open Systems Interconnection—Its Architecture and Protocols. Elsevier Science Publishers, Amsterdam, The Netherlands, 1990. ISBN 0-444-88490-4. xxii + 599 pp. LCCN TK5105.5 .J35 1990.

Jaeschke:MSC89

[Jae89] Rex Jaeschke. Mastering Standard C. Professional Press Books, 101 Witmer Road, Horsham, PA 19044, USA, 1989. ISBN 0-9614729-8-7. LCCN QA76.73.C15 J337 1989. US$39.95. Prepared with L^{A}TEX. Phototypeset using Computer Modern fonts at the American Mathematical Society by arrangement with the publisher.

Jaeschke:DSC91

[Jae91] Rex Jaeschke. The Dictionary of Standard C. Professional Press Books, 101 Witmer Road, Horsham, PA 19044, USA, 1991. ISBN 1-878956-07-8. x + 164 pp. LCCN QA76.73.C15 J335 1991. Prepared with L^{A}TEX, with output to Personal TEX’s 2000dpi phototypesetter using Computer Modern fonts.

Jones:FTV88

[JC88] R. K. Jones and T. Crabtree.

Fortran Tools for VAX/VMS and MS-DOS. John Wiley and Sons, Inc., New York, NY, USA, 1988. ISBN 0-471-61976-0. xii + 447 pp. LCCN QA76.73.F25 J66 1988. Prepared with L^{A}TEX.

JCLT

[JCL89] The Journal of C Language Translation. 1810 Michael Faraday Drive, Suite 101, Reston, VA 22090, USA, Tel: (703) 860-0091.

E-mail rex@aussie.com., 1989.

tj-jcs

[JCo92] Journal of Computer Security, 1992. IOS Press, Amsterdam, The Netherlands. TEX style files are available from the journal editors.

Jen:LCS90-2

[Jen90] Erica Jen. 1989 Lectures in Complex Systems, volume 2. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-50936-9. xxiv + 611 pp. LCCN QA267.7 .C66 1989. Prepared with TEX.

Jones:SSD90

[Jon90] Cliff B. Jones. Systematic Software Development Using VDM. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1990. ISBN 0-13-880733-7. xiv + 333 pp. LCCN QA76.76.D47 J66 1990. Prepared with L^{A}TEX.

Johnson:1998:LAD

[JT98] Michael K. Johnson and Erik W. Troan. Linux Application Development. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-30821-5. xxiii + 538 pp. LCCN QA76.76.O63J635 1998. US$45.95. Prepared with L^{A}TEX.

Stephen:NND90

[Jud90] J. Stephen Judd. Neural Network Design and Complexity of Learning. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-10045-2. 150 pp. LCCN QA76.5 .J83 1990.

Kahlbrandt:SEO98

[Kah98] Bernd Kahlbrandt. Software-Engineering. Objektorientierte Software-Entwicklung mit der Unified Modeling Language. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-63309-X. xvi + 504 pp.

DM 88. URL http://www.
springer.de/catalog/html-
files/deutsch/eng/354063309X.

html. Prepared with L^{A}TEX.

Kamin:PLI90

[Kam90] Samuel N. Kamin. Programming Languages: An Interpreter-Based Approach. Addison-Wes-ley, Reading, MA, USA, 1990. ISBN 0-201-06824-9. xi + 640 pp. LCCN QA76.7.K35 1990. Prepared with L^{A}TEX.

Kane:MEP88

[Kan88] Gordon Kane. Modern Elementary Particle Physics. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-11749-5. xx + 344 pp. LCCN QC793.2 .K361 1987. Prepared with TEX.

Kay:PPP85

[Kay85] David G. Kay. Programming for People/Pascal. Mayfield, Palo Alto, CA, USA, 1985. ISBN 0-87484-717-6. xiii + 620 pp. LCCN QA76.73.P3 K3. Prepared with TEX using Baskerville fonts, with output to Stanford Autologic APS-5 phototypesetter.

Keener:PAM88

[Kee88] James P. Keener. Principles of Applied Mathematics. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-15674-1. xv + 560 pp. LCCN QA601 .K41 1988. Prepared with L^{A}TEX.

Kehoe:ZAI93

[Keh93] Brendan P. Kehoe. Zen and the Art of the Internet: A Beginner’s Guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1993. ISBN 0-13-010778-6. xv + 112 pp. LCCN TK5105.8.I57 K44 1993. Prepared with TEXinfo.

Keller:FCC82

[Kel82] Arthur M. Keller. A First Course in Computer Programming Using PASCAL. McGraw-Hill, New York, NY, USA, 1982. ISBN 0-07-033508-7. xiii + 306 pp. LCCN QA76.73.P2 K44 1982.

Kirkerud:OOP89

[Kir89] Bjørn Kirkerud. Object-Oriented Programming with Simula. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-17574-6. xiii + 515 pp. LCCN QA 76.73 S55 K57 1989. Prepared with L^{A}TEX.

Koonin:CP90

[KM90] Steven E. Koonin and Dawn C. Meredith. Computational Physics. Addison-Wesley, Redwood City, CA, USA, 1990. ISBN 0-201-12779-2. xvi + 635 pp. LCCN QC20.7.E4 K66 1990.

Kreyszig:MCM94

[KN94] E. Kreyszig and E. J. Normington. Maple Computer Manual for Seventh Edition Advanced Engineering Mathematics —Erwin Kreyszig. John Wiley, New York, NY, USA, 1994. ISBN 0-471-31126-X. xii + 506 pp. Prepared with TEX. There is a companion Instructor’s Maple Manual, but I have not yet found an exact citation for it.

Knuth:bible-texts

[Knu91] Donald E. Knuth. 3:16 Bible Texts Illuminated. A-R Editions, Inc., 801 Deming Way, Madison, WI 53717-1903, USA; Tel:

(608) 836-9000, 1991. ISBN 0-89579-252-4. 268 pp. LCCN BS511.2.K58 1991. See also review [She92].

Knuth:AH92

[Knu92] Donald E. Knuth. Axioms and Hulls. Number 606 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992.

ISBN 0-387-55611-7 (New York), 3-540-55611-7 (Berlin). ix + 109 pp. LCCN QA640.3 .K68 1992.

Knuth:1997:FA

[Knu97] Donald E. Knuth. Fundamental Algorithms, volume 1 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, third edition, 1997. ISBN 0-201-89683-4. xix + 650 pp. LCCN QA76.6 .K64 1997. US$49.50.

Knuth:1998:SA

[Knu98a] Donald E. Knuth. Seminumerical Algorithms, volume 2 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, third edition, 1998. ISBN 0-201-89684-2. xiii + 762 pp. LCCN QA76.6 .K64 1997. US$52.75.

Knuth:1998:SS

[Knu98b] Donald E. Knuth. Sorting and Searching, volume 3 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, second edition, 1998. ISBN 0-201-89685-0. xiv + 780 pp. LCCN QA76.6.K64 1997. US$49.95.

Knuth:2000:SPA

[Knu00] Donald E. Knuth. Selected Papers on Analysis of Algorithms. CSLI Publications, Stanford, CA, USA, 2000. ISBN 1-57586-212-3 (paperback), 1-57586-211-5 (cloth). xvi + 621 pp. LCCN QA9.58 .K65 2000.

Knuth:2005:ACPa

[Knu05a] Donald E. Knuth. The Art of Computer Programming: Volume 1, Fascicle 1. MMIX, a RISC Computer for the New Millennium. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-201-85392-2. v + 134 pp. LCCN QA76.6 .K64 2005.

Knuth:2005:ACPb

[Knu05b] Donald E. Knuth. The Art of Computer Programming: Volume 4, Fascicle 3. Generating All Combinations and Partitions. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-201-85394-9. iv + 150 pp. LCCN QA76.6 .K64 2005.

Knuth:2008:ICA

[Knu08] Donald E. Knuth. The Art of Computer Programming: Volume 4, Fascicle 0. Introduction to Combinatorial Algorithms and Boolean Functions. The art of computer programming. Addison-Wesley, Reading, MA, USA, 2008. ISBN 0-321-53496-4. ix + 216 pp. LCCN QA76.6 .K64 2005.

Knuth:2009:ACP

[Knu09] Donald E. Knuth. The Art of Computer Programming: Volume 4, Fascicle 1. Bitwise Tricks and Techniques. Binary Decision Diagrams. The art of computer programming. Addison-Wesley, Reading, MA, USA, 2009. ISBN 0-321-58050-8. vii + 260 pp. LCCN QA76.6 .K64 2009.

Koenderink:SS89

[Koe89] Jan Koenderink. Solid Shape.

The MIT Press, Cambridge, MA, 1989. ISBN 0-262-11139-X. xiv + 699 pp. LCCN QA641 .K74 1990.

Kruse:1999:DSP

[KR99] Robert L. Kruse and Alexander J. Ryba. Data structures and program design in C++. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-768995-0. xvii + 717 pp. LCCN QA76.73.C153 K79 1999. Prepared with TEX. Typeset in two colors with Palatino text and scaled Lucida math fonts.

Krantz:2001:HTM

[Kra01] Steven G. Krantz. Handbook of Typography for the Mathematical Sciences. Chapman and Hall/CRC, Boca Raton, FL, USA, 2001. ISBN 1-58488-149-6. xv + 173 pp. LCCN Z253.4.T47 K65 2001.

Kreps:CMT90

[Kre90] David M. Kreps. A Course in Microeconomic Theory. Princeton University Press, Princeton, NJ, USA, 1990. ISBN 0-691-04264-0. xviii + 850 pp. LCCN HB172 .K74 1990. Typeset using PostScript and Computer Modern fonts.

Kruer:PLP88

[Kru88] William Kruer. The Physics of Laser Plasma Interactions. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-15672-5. xviii + 182 pp. LCCN QC718.5.P5 K781 1988. Prepared with TEX.

Kosekela:PCS91

[KS91] R. Kosekela and M. Simmons, editors. Parallel Computer Systems: Performance Instrumentation and Visualization. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-50937-7. xxii + 289 pp. LCCN T385 .P46 1990. Prepared with L^{A}TEX.

Kassakian:PPE91

[KSV91] John G. Kassakian, Martin F. Shlecht, and George C. Verghese. Principles of Power Electronics. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-09689-7. xxi + 738 pp. LCCN TK7881.15.K37 1991. Prepared with TEX.

Kolb:EU90

[KT90] Edward Kolb and Michael Turner. The Early Universe. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-11603-0. xxii + 547 pp. LCCN QB981 .E68 1990. Prepared with L^{A}TEX.

Kupershmidt:ESS87

[Kup87] Boris A. Kupershmidt. Elements of superintegrable systems: basic techniques and results. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1987. ISBN 90-277-2434-2. xvi + 187 pp. LCCN QA377 .K851 1987. Prepared with TEX.

Lotman:STK92

[L^{+}92] Yu. Lotman et al. Stat’i po tipologii kul’tury. Prilozhenie k ‘Trudam po znakovym sistemam’ (Articles about types of culture. Application to procedures on signed systems.). Tartu Semiotic Group, Tartu, Estonia, 1992. 250 pp.

Langton:AL89

[Lan89] Christopher G. Langton, editor.

Artificial Life. Addison-Wes-ley, Reading, MA, USA, 1989. ISBN 0-201-09346-4 (hardcover), 0-201-09356-1 (paperback). xxix + 655 pp. LCCN QH324.8 .A77 1987. Prepared with TEX.

Levine:MR91

[LB91] R. D. Levine and R. B. Bernstein. Molekulare Reaktionsdynamik. B. G. Teubner, Stuttgart, Germany, 1991. ISBN 3-519-03507-3. Original title: Molecular Reaction Dynamics and Chemical Reactivity, translated by Christian Schlier. The German edition was written in T3, then translated to L^{A}TEX.

Leary:ITF85

[Lea85] Rolfe A. Leary. Interaction Theory in Forest Ecology and Management. Martinus Nijhoff/Dr W. Junk Publishers, Dordrecht, The Netherlands, 1985. ISBN 90-247-3220-4. x + 219 pp. LCCN SD387.M8 L431 1985.

Leslie:RRH91

[Les91] Julia Leslie, editor. Roles and Rituals for Hindu Women. Pinter Publishers, Ltd., 25 Floral St., London WC2E 9DS, UK, 1991. ISBN 0-86187-796-9. xviii + 267 pp. LCCN BL1237.46.R64 1991. Prepared with L^{A}TEX. Used sbTEX 3.1 and Computer Modern fonts, proffed using emTEX’s DVIHPLJ and an HP LaserJet III. Final output produced on bromide at 1270dpi on a Linotronic 300, via DVIPS 5.47 for some parts, and Arbortext’s DVILASER/PS (is that the name?) for others. The L300 is owned and run by the University of London, and they have loaded the CM fonts onto their RIP’s font disk, as well as having the Arbortext software running invisibly, so TeX users can actually submit DVI files over the JANET network (or by direct login and Kermit), and get bromides back through the post 24hrs later.

Laub:MA391

[LHR^{+}91a] Josef Laub, Eugen Hruby, Hans-Christian Reichel, Dieter Litschauer, Herbert Groß, Robert Müller, and Erich Neuwirth. Mathematik Arbeitsbuch 3. Hölder-Pichler-Tempsky Verlag, Vienna, Austria, 1991. ISBN 3-209-01165-6. Textbook for high-school mathematics, grade 7. One full chapter about spreadsheet programs has been typeset by TEX.

Laub:MA491

[LHR^{+}91b] Josef Laub, Eugen Hruby, Hans-Christian Reichel, Dieter Litschauer, Herbert Groß, Robert Müller, and Erich Neuwirth. Mathematik Arbeitsbuch 4. Hölder-Pichler-Tempsky Verlag, Vienna, Austria, 1991. ISBN 3-209-01061-7. Textbook for high-school mathematics, grade 7. One full chapter about spreadsheet programs has been typeset by TEX.

Lim:2010:SHP

[Lim10] An-King Lim. A Sinitic Historical Phonology: Phonological Restructuring of Written Chinese under the 5-th Century Turkic Sinification, volume 40 of Language and Linguistics Monograph Series. Institue of Linguistics, Academica Sinica, Taipei, Taiwan, 2010. ISBN 986-02-5017-0 (hardcover). xvi + 453 pp. LCCN ???? NT$800,

US$60. URL http://www.ling.
sinica.edu.tw/english/v7-2-
review.asp-v_id=48.htm.

Lawrence:NLM87

[LL87] Christopher Lawrence and Ghislaine Lawrence. No Laughing Matter: Historical Aspects of Anaesthesia. Wellcome Institute for the History of Medicine, London, UK, 1987. ISBN 0-85484-056-7. 84 pp. LCCN RD79 .N6 1987. Prepared with L^{A}TEX. PCTEX and Computer Modern fonts, with final output on a Canon LBP8 A1 at 300dpi. The printer was managed by software from Personal TEX driving a Flagstaff Engineering JLaser card plugged into the PC bus, and sending bitmaps directly to the laser printer, bypassing the controller.

Loeckx:FPL88

[LMW88] Jacques Loeckx, Kurt Mehlhorn, and Reinhard Wilhelm. Foundations of Programming Languages. John Wiley, New York, NY, USA, 1988. ISBN 0-471-92139-4. 426 pp. LCCN QA76.7 .L64 1988.

LoPresti:LS91

[LN91] L. Lo Presti and F. Neri. Lanalisi dei segnali. Edizioni Cooperativa Libraria Universitaria Torino, Corso Duca degli Abruzzi, 24, I 10129 Torino, Italy, 1991. Prepared with L^{A}TEX.

Lewis:ETC92

[LP92] H. R. Lewis and Qr’hstos Q. Papadhmhtr’iou (H. H. Papadimitriou). Stoiqe’ia Jewr’ias Ypologismo’u (Elements of the Theory of Computation). Technical Chamber of Greece (“Ekdosh Teqniko’u Epimelhthr’iou ths Ell’adas”), Publications Bureau (“Grafeio Ekdoseon”), c/o Ms. Filippidou, Karageorgi Servias 4, GR-102 48 Athens, Greece, Telephones: +30 - 1 - 32 54 590 (last digit can vary up to … 9), +30 - 1 - 32 45 935, Telefax: +30 - 1 - 32 21 772, 1992. 494, with subject index, author index, and glossary pp. One of the first books to be typeset in Greek by TEX.

Lyche:1989:MMC

[LS89] Tom Lyche and Larry L. Schumaker, editors. Mathematical Methods in Computer Aided Geometric Design. Academic Press, New York, NY, USA, 1989. ISBN 0-12-460515-X. LCCN QA448.D38 M381 1989.

Lusardi:DEG88

[Lus88] Frank Lusardi. The Database Experts’ Guide to SQL. McGraw-Hill, New York, NY, USA, 1988. ISBN 0-07-039006-1 (hardcover), 0-07-039002-9 (paperback). xiv + 224 pp. LCCN QA76.73.S67 L87 1988. Prepared with TEX.

Lynch:1996:DA

[Lyn96] Nancy A. Lynch. Distributed Algorithms. Morgan Kaufmann series in data management systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1996. ISBN 1-55860-348-4. xvii + 872 (or xxiii + 872??) pp. LCCN QA76.9.D5 L96 1996. Prepared with L^{A}TEX.

Maeder:PM90

[Mae90] Roman Maeder. Programming in Mathematica. Addison-Wes-ley, Reading, MA, USA, 1990. ISBN 0-201-51002-2. xiv + 267 pp. LCCN QA76.95 .M34 1989. Prepared with TEX.

MathSoft:1985:PUG

[Mat85] MathSoft, Inc. S-PLUS User’s Guide: Version 3.3 for Windows. StatSci, a Division of MathSoft, Inc., Seattle, WA, USA, September 1985. various pp. Prepared with L^{A}TEX.

Murray:EMI90

[MB90] David Murray and Bernard Buxton. Experiments in the Machine Interpretation of Visual Motion. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-13263-X. 236 pp. LCCN TA1635 .M87 1990.

Mangel:DMB88

[MC88] Marc Mangel and Colin W. Clark. Dynamic Modeling in Behavioral Ecology. Princeton University Press, Princeton, NJ, USA, 1988. ISBN 0-691-08505-6 (hardcover), 0-691-08506-4 (paperback). xii + 308 pp. LCCN QL751.65.M3 M291 1988.

Misner:SP91

[MC91] Charles W. Misner and Patrick J. Cooney. Spreadsheet Physics. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-16410-8. xvii + 228 pp. LCCN QC52 .M57 1991. Prepared with L^{A}TEX.

McAllester:O89

[McA89] David McAllester. Ontic. The MIT Press, Cambridge, MA, 1989. ISBN 0-262-13235-4. xiv + 158 pp. LCCN QA76.9.A96 M37 1989.

McCarthy:ITK90

[McC90] J. M. McCarthy. Introduction to Theoretical Kinematics. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-13252-4. 130 pp. LCCN QA841 .M33 1990.

Mead:AVL89

[Mea89] Carver Mead. Analog VLSI and Neural Systems. Addison-Wes-ley, Reading, MA, USA, 1989. ISBN 0-201-05992-4. xxii + 371 pp. LCCN QA76.5 .M48 1989. Prepared with TEX.

Mesterton-Gibbons:CAM89

[MG89] Michael Mesterton-Gibbons. A Concrete Approach to Mathematical Modeling. Addison-Wes-ley, Reading, MA, USA, 1989. ISBN 0-201-12910-8. xx + 597 pp. LCCN QA401 .M51381 1989. Prepared with TEX.

Milanese:1996:CIL

[Mil96] Guido Milanese. Concordantia et instrumenta lexicographica ad Graduale Romanum pertinentia. Editrice Liguria s.n.c. di Norberto Sabatelli & C., Via De Mari 4R - I-17100 Savona, Italy; or Postal Box 181 - 17100 Savona, Italy, 1996. ISBN 88-8055-151-5. 550 pp. US$55.00 (80,000 Italian Lire). Preface by Mirella Ferrari. The book is a concordance and a collection of indexes done on the Graduale Romanum, the most important source of the Latin liturgy. The typesetting of the book is completely done with L^{A}TEX. There are several kinds of indexes, and — due to the variety of such indexes — several kinds of L^{A}TEX packages were used, including multicols, various kinds of lists and a huge use of the myverb package (hundreds of pages!), because I wanted to produce the concordance (not the indexes) at fixed space. Copies for review on journals available on request:

Contact the author at mc1194@

mclink.it.

Maslen:PAW91

[ML91] Keith Maslen and John Lancaster, editors. The Bowyer Ledgers: The printing accounts of William Bowyer father and son reproduced on microfiche with a checklist of Bowyer printing 1699-1777, a commentary, indexes, and appendixes. The Bibliographical Society and The Bibliographical Society of America, London, UK and New York, NY, USA, 1991. ISBN 0-914930-13-3. 70 microfiches pp. LCCN Z232.B68.B6 1991. The editor (JL) remarks: “The printed volume contains lxxv, 616 pp., being the apparatus and indexes to 70 microfiche which reproduce the original documents. It may be of interest to note that the estimate for commercial typesetting (from existing well-marked ASCII files, but involving many complications such as tables) was in excess of $20,000; the total cost for TEX was below $2,000 (plus of course one editor’s unpaid time – but for this sort of work, with a very small market, the key is getting it published at all). In addition, there was the great advantage that finetuning could take place right up to the last minute, with instant proofs at low cost. I am (as a former commercial printer as well as a printing historian and hobbyist letterpress printer) thoroughly sold on TEX, and am using it for three additional books now in progress – for technical bibliography, which can often be as complicated as mathematics, the range of possibilities and flexibility makes it well worth the complexity.”.

Moye:1995:FTD

[Moy95] Stephen Moye. Fontographer:

Type by Design. MIS:Press, New York, NY, USA, 1995. ISBN 1-55828-447-8. xi + 275 pp. LCCN Z250.8.F65 M69 1995. US$30.00 (est.). Prepared with TEX.

Maurer:DAM91

[MR91] Stephen B. Maurer and Anthony Ralston. Discrete Algorithmic Mathematics. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-15585-0. xix + 834 pp. LCCN QA39.2 .M394 1991. Prepared with TEX.

Mason:RHM85

[MS85] M. T. Mason and J. K. Salisbury, Jr. Robot Hands and the Mechanics of Manipulation. The MIT Press, Cambridge, MA, 1985. ISBN 0-262-13205-2. xxvi + 298 pp. LCCN TJ211 .M366 1985. Prepared with TEX (unknown variety).

Morner:CSE88

[MS88] Magnus Möner and Thommy Svensson, editors. Classes Strata Elites: Essays on Social Stratification in History. Number 34 in Report from the Department of History, Gothenburg University. Acta Universitatis Gothoburgensis, Gothenburg, Sweden, 1988.

ISBN 91-7970-241-4. ISSN 0346-914X. 269 pp. LCCN HN13 .C53 1988.

Munkres:AM91

[Mun91] James Munkres. Analysis of Manifolds. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-51035-9. xi + 366 pp. LCCN QA300 .M75 1991.

Manna:LBC90-1

[MW90a] Zohar Manna and Richard Waldinger. The Logical Basis for Computer Programming, volume 1. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-18260-2. LCCN QA76.6 .M35595 1985. Prepared with TEX.

Manna:LBC90-2

[MW90b] Zohar Manna and Richard Waldinger. The Logical Basis for Computer Programming, volume 2. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-18261-0. LCCN QA76.6 .M35595 1985. Prepared with TEX.

MacCallum:ACR91

[MW91] Malcolm A. H. MacCallum and Francis J. Wright, editors. Algebraic Computing with REDUCE —Lecture Notes from the First Brazilian School on Computer Algebra. Clarendon Press, Oxford, UK, 1991. ISBN 0-19-853443-4 (paperback), 0-19-853444-2 (hardcover). LCCN QA155.7.E4 B73 1989. US$29.95 (paperback), US$45.00 (hardcover). Prepared with L^{A}TEX.

Nakamura:TRR91

[Nak91] Yoshihiko Nakamura. Theoretical Robotics: Redundancy and Optimization. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-15198-7. xi + 337 pp. LCCN TJ211 .N34 1991. Prepared with TEX.

Nash:HSC90

[Nas90] Stephen G. Nash, editor. A History of Scientific Computing. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-50814-1. xix + 359 pp. LCCN QA76.17 .H59 1990. Prepared with TEX.

Neural-Computation

[NC89] Neural Computation. The MIT Press, 1989. ISSN 0899-7667. Technical journal published four times per year, using Post-Script and Computer Modern fonts.

Newton:TTS88

[New88] H. Joseph Newton. Timeslab: A Time Series Analysis Laboratory. Wadsworth and Brooks/Cole Publishing Company, Pacific Grove, CA, USA, 1988. ISBN 0-534-09198-9. xv + 623 pp. LCCN QA280 .N487 1988. Prepared with TEX.

Nickson:HDP93

[Nic93] R. Andrew Nickson. Historical dictionary of Paraguay. Scarecrow Press, Metuchen, NJ, USA, second edition, 1993. ISBN 0-8108-2643-7. xxi + 685 pp. LCCN F2664 .K64 1993. Typeset with plain TEX using emtex on an IBM PC.

Nossal:ICB90

[NL90] Ralph Nossal and Harold Lecar. Introduction to Cell Biophysics. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-19560-7. xii + 387 pp. LCCN QH581.5 .N67 1989. Prepared with TEX.

Newell:NLO91

[NM91] Alan C. Newell and Jerome V. Molony. Non-Linear Optics. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-51014-6. xi + 436 pp. LCCN QC446.2.N48 1992.

Negele:QMP88

[NO88] John W. Negele and Henri Orland. Quantum Many-Particle Systems. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-12593-5. xviii + 459 pp. LCCN QC174.17.P7 N441 1988. Prepared with TEX.

Othmer:1997:CSM

[OALD97] H. G. Othmer, F. R. Adler, M. A. Lewis, and J. C. Dallon, editors. Case Studies in Mathematical Modeling—Ecology, Physiology, and Cell Biology. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-574039-8. viii + 411 pp. LCCN QH541.15.M3C37

1997.  US$41.00.  URL

https://www.math.utah.edu/
books/csmmepcb/index.html;
 http://www.phptr.com/. Pre-

pared with L^{A}TEX.

Orlando:FAS91

[OD91] Terry P. Orlando and Kevin A.

Delin. Foundations of Applied Superconductivity. Addison-Wes-ley, Reading, MA, USA, 1991. ISBN 0-201-18323-4. xx + 584 pp. LCCN QC611.92 .O75 1991. Prepared with L^{A}TEX.

OKeefe:CP90

[O’K90] Richard O’Keefe. The Craft of Prolog. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-15039-5. xix + 387 pp. LCCN QA76.73.P76 O38 1990.

Omondi:CAS94

[Omo94]  Amos R. Omondi.  Com-

puter Arithmetic Systems— Algorithms, Architecture, Implementation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-334301-4. xvi + 520 pp. LCCN QA76.9.C62 O46 1994. US$40.00. Prepared with L^{A}TEX, using a style file provided by the publisher. PostScript page proofs were prepared by the author, and delivered to the publisher for the final printing.

Orlando:LDA88

[OP88] M. Orlando and G. Podda. Lineamenti di disegno automatico. Edizioni Cooperativa Libraria Universitaria Torino, Corso Duca degli Abruzzi, 24, I 10129 Torino, Italy, 1988. Prepared with TEX by Tecnograf.

OShaughnessy:SC87

[O’S87] Douglas O’Shaughnessy. Speech Communications. Addison-Wes-ley, Reading, MA, USA, 1987. ISBN 0-201-16520-1. xviii + 568 pp. LCCN P95.O74 1987. Prepared with TEX.

Palmer:SCW89

[Pal89] Richard Palmer. A Supplementary Catalogue of Western Manuscripts in the Library of the Wellcome Institute for the History of Medicine: First Cumulative Supplement. Wellcome Institute for the History of Medicine, London, UK, 1989. Prepared with L^{A}TEX. PCTEX and Computer Modern fonts, with final output on an HP LaserJet II at 300dpi.

Pao:APR89

[Pao89] Yoh-Han Pao. Adaptive Pattern Recognition and Neural Networks. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-12584-6. xviii + 309 pp. LCCN TK7882.P3 P36 1989. Prepared with TEX.

Parks:ESP91

[Par91] George Parks. Electrodynamics of Space Plasmas. Addison-Wes-ley, Reading, MA, USA, 1991. ISBN 0-201-50821-4. ix + 538 pp. LCCN QC809.P5 P37 1991. Prepared with L^{A}TEX.

Pardalos:CNO93

[Par93] Panos M. Pardalos, editor. Complexity in Numerical Optimization. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-1415-4. xiii + 511 pp. LCCN QA 267.7 P37 1993. Prepared with TEX.

Paul:RMM81

[Pau81] Richard P. Paul. Robot Manipulators: Mathematics, Programming and Control. The MIT Press, Cambridge, MA, 1981. ISBN 0-262-16082-X. 279 pp. LCCN TS191.8 .P38. Another early book set in plain TEX.

Pavone:IIF92

[Pav92] Marco Pavone. Integrali impropri e funzioni integrali (Improper integrals and integral functions): Esercizi e complementi di Analisi Matematica I (English: Exercises and complements to Mathematical Analysis I). Aracne Editrice, Via Luigi Chiozza, 21b, 00133 Roma, Italy, 1992. ISBN 88-7999-000-4. Prepared with L^{A}TEX. The final formatting was done by TEXmat, Via Luigi Chiozza, 31a/b, 00133 Roma, Italy.

Pavone:TES93

[Pav93] Marco Pavone. Temi di esame svolti di analisi amtematica I (English: Solved examination problems in mathematical analysis I). Aracne Editrice, Via Luigi Chiozza, 21b, 00133 Roma, Italy, 1993. ISBN 88-7999-002-0. Prepared with L^{A}TEX. The final formatting was done by TEXmat, Via Luigi Chiozza, 31a/b, 00133 Roma, Italy.

Parker:PNA89

[PC89] T. S. Parker and L. O. Chua. Practical Numerical Algorithms for Chaotic Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-96689-7 (USA) and 3-540-96689-7 (Germany). xiv + 348 pp. LCCN Q172.5.C45 P37 1989. Prepared with L^{A}TEX.

PDG:PPDB90

[PDG90] Particle properties data booklet. Amsterdam, The Netherlands, April 1990. Reprinted from Physics Letters B239.

Perelson:TI88-1

[Per88a] Alan S. Perelson. Theoretical Immunology, volume 1. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-15688-1. LCCN QR182.2.M36 T485 1987. Prepared with TEX.

Perelson:TI88-2

[Per88b] Alan S. Perelson. Theoretical Immunology, volume 2. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-15683-0. LCCN QR182.2.M36 T485 1987. Prepared with TEX.

Pertritskijj:VK92

[Pet92a] Velemir Petritskijj. Veka krepostnojj (English: Serf of the Century). publisher unknown, address unknown, 1992. 54 pp.

Petrov:RA92

[Pet92b]  Viktor Petrov.  Russkie v

Amerike: 20 Vek. (Russians in America: 20th Century). Russian-American Historical Society, Washington, DC, USA, 1992. ISBN 0-911971-67-X. 149 pp. LCCN E184.R9 P475 1992.

Press:NRA86

[PFTV86] William H. Press, Brian P. Flannery, Saul A. Teukolsky, and William T. Vetterling. Numerical Recipes—The Art of Scientific Computing. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-30811-9. xx + 818 pp. LCCN QA297 .N866 1986. Originally for FORTRAN and then rewritten for a C version, Numerical Recipes in C ; also example books in Fortran, Pascal and C, making a total of five books with “Typeset in TEX” on the back of the title page.

Piccolo:TBT92

[Pic92] E. Piccolo. Tecniche di base e tecniche avanzate per l’uso del personal computer (Base and advanced techniques for using personal computers). Edizioni Cooperativa Libraria Universitaria Torino, Corso Duca degli Abruzzi, 24, I 10129 Torino, Italy, 1992. Set in LaTeX with a customized style named clutful as mentioned in the back of the titlepage. Uses the 12pt option with standard Computer Modern type faces. The illustrations are mainly hard copies of computer screens reduced and pasted in place. \tt type face is used a lot for reproducing computer programs.

Pierce:BCT91

[Pie91] Benjamin Pierce. Basic Category Theory for Computer Scientists. The MIT Press, Cambridge, MA, 1991. ISBN 0-262-66071-7. xiii + 100 pp. LCCN QA76.9.M35 P54 1991. Prepared with L^{A}TEX.

Piff:DM91

[Pif91] Mike Piff. Discrete Mathematics —An introduction for software engineers. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-38475-3 (hardcover), 0-521-38622-5 (paperback). xi + 317 pp. LCCN QA76.9.M35 P544 1991. Prepared with L^{A}TEX, with some help from plain TEX.

Peitgen:CF92

[PJS92] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Chaos and Fractals: New Frontiers of Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97903-4, 3-540-97903-4. xvi + 984, with 686 illustrations, 40 in color pp. LCCN QA614.86 .P43 1992. US$49. Prepared with L^{A}TEX.

Peitgen:CF93

[PJS93] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Chaos and Fractals: New Frontiers of Science (corrected second printing). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97903-4, 3-540-97903-4. xvi + 984, with 686 illustrations, 40 in color pp. LCCN QA614.86 .P43 1993. US$49. Prepared with L^{A}TEX.

Perelson:MER91-9

[PK91] Alan S. Perelson and Stuart A.

Kauffman. Molecular Evolution on Rugged Landscapes: Proteins, RNA and the Immune System, volume 9. Addison-Wes-ley, Reading, MA, USA, 1991. ISBN 0-201-52149-0 (hardcover), 0-201-52150-4 (paperback). xviii + 309 pp. LCCN QH325 .W67 1989. Prepared with TEX.

Pozar:ME90

[Poz90] David M. Pozar. Microwave Engineering. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-50418-9. xvii + 726 pp. LCCN TK7876 .P69 1990. Prepared with TEX.

AFIT:CTN91

[Prz91] J. S. Przemieniecki, editor. Air Force Institute of Technology —Critical Technologies for National Defense. American Institute of Aeronautics and Astronautics, 370 L’Enfant Promenade SW, Washington, DC 20024–2518, 1991. ISBN 1-56347-009-8. 325 pp. LCCN UG633 .C74 1991. Prepared with L^{A}TEX.

Pieprzyk:DHA93

[PS93] Josef Pieprzyk and Babak Sadeghiyan. Design of Hashing Algorithms. LNCS 756. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-57500-6. xiii + 194 pp. LCCN QA76.9.H36 P53 1993. Prepared with L^{A}TEX.

Piatetski:KDD91

[PSF91] Gregory Piatetski-Shapiro and William Frawley, editors. Knowledge Discovery in Databases. The MIT Press, Cambridge, MA, 1991. ISBN 0-262-66070-9 (paper). xii + 525 pp. LCCN Q325.5 .K68 1991. Prepared with L^{A}TEX.

Rahtz:info-technology

[Rah87] Sebastian P. Q. Rahtz, editor. Information Technology in the Humanities, Computers and their applications. Ellis Horwood, New York, NY, USA, 1987. ISBN 0-7458-0148-X. LCCN AZ105 .I541 1987. UK£25.00. Laser-printed using L^{A}TEX.

Raibert:LR86

[Rai86] Mark X. Raibert. Legged Robots That Balance. The MIT Press, Cambridge, MA, 1986. ISBN 0-262-18117-7. xiii + 233 pp. LCCN TJ211 .R351 1986. Prepared with TEX.

Ramond:FTM90

[Ram90] Pierre Ramond. Field Theory: A Modern Primer. Addison-Wesley, Reading, MA, USA, second edition, 1990. ISBN 0-201-54611-6. xix + 329 pp. LCCN QC174.45 .R35 1990. Prepared with TEX.

Raman:1998:AST

[Ram98] T. V. Raman. Audio system for technical readings, volume 1410 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-65515-8. xvi + 121 pp. LCCN QA267.A1 L43 no.1410. Originally presented as the author’s thesis (Ph. D.– Cornell University, 1994).

Rashid:CMU91

[Ras91] Richard Rashid, editor. CMU Computer Science; A 25th Anniversary Commemorative. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-52899-1. xxxii + 519 pp. LCCN QA75.5 .C548 1990. Prepared with L^{A}TEX.

Raudsepp:GK92

[Rau92] Jaanus Raudsepp. Gogol’ v KGB: a documentary novel. Effect Publishing, 501 Fifth Avenue, Suite 1612, New York, NY 10017, USA, 1992. ISBN 0-911971-72-6. 141 pp. LCCN PG3485.5.A93 G64 1992.

Robertson:EL91

[RBM^{+}91] David Robertson, Alan Bundy, Robert Muetzelfeldt, Mandy Haggith, and Michael Uschold. Eco-Logic. The MIT Press, Cambridge, MA, 1991. ISBN 0-262-18143-6. 243 pp. LCCN QH541.15.S5 E26 1991. Prepared with L^{A}TEX.

Ricco:ICI91

[RFB91] B. Riccó, F. Fantini, and P. Brambilla. Introduzione ai circuiti integrati digitali. Zanichelli-Telettra, Bologna, Italy, 1991. ISBN 88-08-07838-8. viii + 445 pp. The book has been composed in L^{A}TEX with minor changes to the standard style ‘book.sty’, specifically with new headers and new ‘empty’ pagestyle for typesetting the initial page of each chapter. The book has been set with the 12pt option, laserprinted with a resolution of 300 dpi, and reproduced with no reduction. The characters are the standard ones of the Computern Modern series. The whole work was done by the authors, and the publisher limited himself to reproducing the laser-printed sheets received from the authors, and to preparing the titlepage and the cover.

Richards:CFL89

[Ric89] Tom Richards. Clausal Form Logic. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-12920-5. xiii + 359 pp. LCCN QA76.76.E95 R531 1989. Prepared with TEX.

Robertson:1995:EMM

[Rob95] John S. Robertson. Engineering mathematics with Mathematica. International series in pure and applied mathematics. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-053171-4. xiv + 288 pp. LCCN TA345.R55 1995. Prepared with L^{A}TEX.

Rogers:PEC85

[Rog85] David F. Rogers. Procedural Elements for Computer Graphics. McGraw-Hill, New York, NY, USA, 1985. ISBN 0-07-053534-5. xiii + 433 pp. LCCN T385 .R631 1985. US$19.95.

Rosbach:U90

[Ros90] Johan Hammond Rosbach. Unumane (Single-handedly). Esperantoforlaget AS, Oslo, Norway, 1990. ISBN 82-90006-09-8. Prepared with L^{A}TEX. This is a novel in Esperanto.

Rose:SB91

[Ros91] Marshall T. Rose. The Simple Book. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-812611-9. xxix + 347 pp. LCCN TK5105.5.R68 1991.

Robbins:CXM89

[RR89] Kay A. Robbins and Steven Robbins. The Cray X-MP/Model 24. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-97089-4, 3-540-97089-4. vi + 165 pp. LCCN QA76.8 C72 R63 1989. Prepared with L^{A}TEX.

Reilly:AIA92

[RR92] Paul Reilly and Sebastian Rahtz.

Archaeology and the Information Age. Routledge, London, UK, 1992. ISBN 0-415-07858-X. xxiv + 395 pp. LCCN CC80.4 .A73 1992. Prepared with L^{A}TEX.

Richardson:TCM88

[RS88] Robert C. Richardson and Eric N. Smith. Techniques in Condensed Matter: Physics at Low Temperatures. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-15002-6. xiii + 338 pp. LCCN QC173.4.C65 E97 1988. Prepared with TEX.

Rich:PA90

[RW90] Charles Rich and Richard C. Waters. The Programmer’s Apprentice. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-52425-2. xi + 238 pp. LCCN QA76.76.D47 R53 1990. Prepared with L^{A}TEX.

Raymond:NHD91

[Rw91] Eric Raymond and with assistance and illustrations by Guy L. Steele Jr., editors. The New Hacker’s Dictionary. The MIT Press, Cambridge, MA, 1991. ISBN 0-262-18145-2 (hardcover), 0-262-68069-6 (paperback). xx + 433 pp. LCCN PN6231.E4 H3 1991. This book was typeset by E.S.R. using TEX 3.0 and a customized version of the TEXinfo package; eleventh-hour assistance was provided by Mark Hoover at Digital Graphics, Inc. DVI-to-PS conversion was done using Y&Y’s DVIPSONE. The camera-ready copy was generated off a L300 phototypesetter by Chiron, Inc. The running text was set in Adobe’s version of Palatino. This book corresponds to version 2.9.6 of the on-line Jargon File.

Sabot:PM88

[Sab88] Gary Sabot. The Paralation Model. The MIT Press, Cambridge, MA, 1988. ISBN 0-262-19277-2. x + 238 pp. LCCN QA76.6 .S2161 1988.

Sadiku:NTE92

[Sad92] Matthew N. O. Sadiku. Numerical Techniques in Electromagnetics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1992. ISBN 0-8493-4232-5. xviii + 690 pp. LCCN QC760 .S24 1992. Prepared with AMS-TEX.

Salomon:1995:AT

[Sal95]  David Salomon.  The Ad-

vanced TEXbook. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94556-3. xx + 490 pp. LCCN Z253.4.T47 S25 1995. US$39.95. URL

http://www.ecs.csun.edu/~
dxs/tatb.advertis/tatbAd.
html;   http://www.booksbydavidsalomon.
com/.

Salomon:1998:DCC

[Sal98] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98280-9. xx + 427 pp. LCCN QA76.9.D33S25

1997.  US$39.95.  URL

http://www.ecs.csun.edu/~
dxs/DCadvertis/DcompAd.html;
 http://www.booksbydavidsalomon.
com/.

Salomon:1999:CGG

[Sal99]  David Salomon.  Computer

Graphics and Geometric Modeling. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98682-0. xviii + 851 pp. LCCN T385 S243 1999. URL

                                        http://www.ecs.csun.edu/~
                                        dxs/CGGMadvertis/CGad.html;
                                         http://www.booksbydavidsalomon.
                                        com/.

Salomon:2000:DCC

[Sal00] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 0-387-95045-1. xvi + 823 pp. LCCN QA76.9.D33 S25

2000.  US$39.95.  URL

                                        http://www.ecs.csun.edu/~
                                        dxs/DCadvertis/Dcomp2Ad.html;
                                         http://www.booksbydavidsalomon.
                                        com/.

Salomon:2001:GDC

[Sal01]  David Salomon.  A guide

to data compression methods. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-95260-8. xii + 295 pp. LCCN QA76.9.D33

                                        S28 2001. URL http://www.
                                        booksbydavidsalomon.com/.

Salomon:2003:DPS

[Sal03]  David Salomon.  Data Pri-

vacy and Security. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-00311-8. xiv + 465 pp. LCCN QA76.9.A25 S265 2003. US$59.95. URL http://www.

booksbydavidsalomon.com/.

Salomon:2004:DCC

[Sal04] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2004. ISBN 0-387-40697-2. xx + 898 pp. LCCN QA76.9.D33 S25 2004. URL http://

www.ecs.csun.edu/~dsalomon/
DC3advertis/DComp3Ad.html;
http://www.booksbydavidsalomon.
com/.

Salomon:2005:CDC

[Sal05a] David Salomon. Coding for data and computer communications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-21245-0. xv + 548 pp. LCCN TK5102.94 .S35 2005. URL http://www.

DavidSalomon.name/Codes/Codes.
html;   http://www.ecs.csun.
edu/~dsalomon/.

Salomon:2005:FCS

[Sal05b] David Salomon. Foundations of computer security. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-84628-193-8. xxi + 368 pp. LCCN ???? URL http:/

/www.springer.com/sgw/cda/

frontpage/0,11855,4-40007-
22-65173048-0,00.html;   http:
//www.springer.com/sgw/cda/
pageitems/document/cda_downloaddocument/
0,11855,0-0-45-166689-p65173048,
00.pdf;   http://www.springer.
com/sgw/cda/pageitems/document/
cda_downloaddocument/0,11855,
0-0-45-166687-p65173048,00.
pdf;   http://www.springer.
com/sgw/cda/pageitems/document/
cda_downloaddocument/0,11855,
0-0-45-166688-p65173048,00.
pdf.

Salomon:2006:CSC

[Sal06a] David Salomon. Curves and Surfaces for Computer Graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-24196-5, 0-387-28452-4 (ebook). xvi + 460 pp. LCCN T385 .S2434 2005.

Salomon:2006:TPC

[Sal06b] David Salomon. Transformations and Projections in Computer Graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-84628-392-2. xiv + 288 pp. LCCN T385 .S2439 2006. US$59.95.

Salomon:2007:DCC

[Sal07] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-84628-602-6. xxv + 1092 pp. LCCN ???? With contributions by Giovanni Motta and David Bryant.

Scharf:FCE90

[SB90] Louis L. Scharf and Richard T.

Behrens. A First Course in Electrical and Computer Engineering with MATLAB. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-53472-X. xv + 269 pp. LCCN TK168 .S34 1990. Prepared with TEX.

Simmons:IFP89

[SBK89] Margaret Simmons, Ingrid Bucher, and Rebecca Koskela. Instrumentation for Future Parallel Computer Systems. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-50390-5. xx + 260 pp. LCCN QA76.5 .I48751 1989. Prepared with L^{A}TEX.

Schmueser:FGE91

[Sch88]  Peter Schmüser.  Feynman-

Graphen und Eichtheorien für Experimentalphysiker. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 3-540-18797-9. vi + 217 pp. LCCN QC794.6.F4 S35 1988.

Scharf:SSP91

[Sch91] Louis L. Scharf. Statistical Signal Processing: Detection, Estimation, and Time Series Analysis. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-19038-9. xviii + 524 pp. LCCN TK5102.5 .S3528 1991. Prepared with TEX.

Sedgewick:A83

[Sed83] Robert Sedgewick. Algorithms. Addison-Wesley, Reading, MA, USA, 1983. ISBN 0-201-06672-6. viii + 551 pp. LCCN QA76.6 .S435 1983.

Sedgewick:1988:A

[Sed88] Robert Sedgewick. Algorithms. Addison-Wesley, Reading, MA, USA, second edition, 1988. ISBN 0-201-06673-4. xii + 657 pp. LCCN QA76.6 .S435 1988. US$34.95.

Sedgewick:AC90

[Sed90] Robert Sedgewick. Algorithms in C. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-51425-7. xii + 657 pp. LCCN QA76.73.C15 S43 1990. Prepared with TEX.

Sehlinger:1992:UGV

[Seh92] Bob Sehlinger. The unofficial guide to Las Vegas. Prentice-Hall Travel, New York, NY, USA, 1992. ISBN 0-13-950833-3. xii + 495 pp. LCCN F849.L35 S44 1992. US$12.00. Prepared with TEX. Typesetting by Tseng Information Systems, 621 William Vickers Avenue, Durham, NC 27701, USA.

Springer:SAP89

[SF89] George Springer and Daniel Friedman. Scheme and the Art of Programming. The MIT Press, Cambridge, MA, 1989. ISBN 0-262-19288-8. xxiv + 595 pp. LCCN QA76.6 .S6861 1989.

Sharp:SCA90

[Sha90] R. Y. Sharp. Steps in Commutative Algebra. London Mathematical Society Student Texts 19. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-39338-8 (hardcover), 0-521-39732-4 (paperback). xi + 321 pp. LCCN QA251.3 .S53 1990. Prepared with L^{A}TEX.

Shein:1992:BRB

[She92]  Barry Shein.  Book review:

3:16 Bible Texts Illuminated. ;login: the USENIX Association newsletter, 17(1):31–??, January/February 1992. CODEN LOGNEM. ISSN 1044-6397.

Sharples:CT89

[SHH^{+}89] Mike Sharples, David Hogg, Chris Hutchison, Steve Torrance, and David Young. Computers and Thought. The MIT Press, Cambridge, MA, 1989. ISBN 0-262-19285-3. xxix + 401 pp. LCCN Q335 .C57 1989.

Siebert:CSS85

[Sie86] William McC. Siebert. Circuits, Signals, and Systems. The MIT Press and McGraw-Hill, Cambridge, MA, 1986. ISBN 0-262-19229-2 (MIT Press) and 0-07-057290-9 (McGraw-Hill). xvi + 651 pp. LCCN TK454 .S57 1986. Prepared with TEX.

Sigler:2002:FLA

[Sig02] L. E. (Laurence E.) Sigler. Fibonacci’s Liber Abaci: A Translation into Modern English of Leonardo Pisano’s Book of Calculation. Sources and studies in the history of mathematics and physical sciences. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95419-8. viii + 636 pp. LCCN QA32 .F4713 2002. US$99.00. Prepared with TEX.

Siemens:ENM87

[SJ87] Phillip J. Siemens and Aksel S. Jensen. Elements of the Nuclei: Many-Body Physics with the Strong Interaction. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-15572-9. xiii + 369 pp. LCCN QC793.3.S8 S541 1987. Prepared with TEX.

Sexl:P91

[SKPS91] Roman Sexl, Helmut Kuehnelt, Alfred Pflug, and Helga Stadler. Physik 3. Hölder-Pichler-Tempsky Verlag, Vienna, Austria, 1991. ISBN 3-209-01191-5. Textbook for high-school physics, grade 7. The whole book has been typeset by TEX.

Smith:IAP90

[Smi90] Stanley H. Smith. Investigations of the Attack on Pearl Harbor: Index to Government Hearings, volume 3 of Bibliographies and Indexes In Military Studies. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1990. ISBN 0-313-26884-3. ix + 250 pp. LCCN D767.92 .I58 1990.

Sanchez-Mondragon:LMO86

[SMW86] J. Sánchez-Mondragón and K. B.

Wolf, editors. Lie Methods in Optics, volume 250 of Lecture Notes in Physics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-16471-5. xiv + 247 pp. LCCN QC355.2 .L54 1985. DM49.00.

Sokolsky:IUE89

[Sok89] Pierre V. Sokolsky. Introduction to Ultrahigh Energy Cosmic Ray Physics. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-17634-3. xvi + 208 pp. LCCN QC485 .S53 1988. Prepared with L^{A}TEX.

tj-solstice

[Sol92] Solstice: An Electronic Journal of Geography and Mathematics, 1992. ISBN 1-877751-44-8; 1-877751-52-9; 1-877751-53-7. ISSN 1059-5325. Institute of Mathematical Geography, 2790 Briarcliff, Ann Arbor, MI 48105-1429, USA.

Sosinksij:KNM91

[Sos92]  A. Sosinksij.  Kak napisat’

matematicheskuyu stat’yu poanglijski (How to prepare English-language mathematical papers). Sophus Lie Mathematical Center, Minsk, Byelorussia, 1992.

Spector:ISS94

[Spe94] Phil Spector. An Introduction to S and S-Plus. Duxbury Press, Belmont, CA, USA, 1994. ISBN 0-534-19866-X. xiii + 286 pp. LCCN QA76.73.S15 S63 1994. US$28.95.

Sterling:AP90

[SS86] Leon Sterling and Ehud Shapiro. The Art of Prolog. The MIT Press, Cambridge, MA, 1986. ISBN 0-262-19250-0 (hardcover), 0-262-69105-1 (paperback). xxviii + 437 pp. LCCN QA76.73.P76 S74 1986.

Scheck:MM91

[SS89] Florian Scheck and Rainer Schöpf. Mechanik-Manual. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51211-X. Prepared with L^{A}TEX. Final copy prepared on a 300dpi printer, then photoreduced.

Sullivan:AII91

[ST91] Joseph W. Sullivan and Sherman W. Tyler, editors. Architecture for Intelligent Interfaces. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-50305-0. xv + 472 pp. LCCN QA76.9.H85 I583 1991. Prepared with L^{A}TEX.

Steele:1990:CLL

[Ste90a] Guy L. Steele Jr. Common Lisp: The Language. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, second edition, 1990. ISBN 1-55558-041-6 (paperback), 1-55558-042-4 (hardcover), 0-08-050226-1 (e-book), 0-13-152414-3 (Prentice-Hall). xxiii + 1029 pp.

LCCN QA76.73.L23 S73 1990. With contributions by Scott E. Fahlman and others, and with contributions to the second edition by Daniel G. Bobrow and others. See also [?, ?].

Sterling:PP90

[Ste90b] Leon Sterling, editor. The Practice of Prolog. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-19301-9. 312 pp. LCCN QA76.73.P76 P73 1990.

Stevick:FOE92

[Ste92] Robert D. Stevick. A Firstbook of Old English. Kenkyusha, Tokyo, Japan, 1992. ISBN 4-327-40104-8.

Strack:GM89

[Str89] Otto D. L. Strack. Groundwater Mechanics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-365412-5. xix + 732 pp. LCCN TC176 .S77 1989. Prepared with TEX.

Strogat:BLT92

[Str92] Lew Strogat. Both Laughter and Tears: a Collection of Stories. Effect Publishing, 501 Fifth Avenue, Suite 1612, New York, NY 10017, USA, 1992. ISBN 0-911971-69-6.

Stutz:2001:LCT

[Stu01] Michael Stutz. The Linux Cookbook: Tips and Techniques for Everyday Use. Linux Journal Press, San Francisco, CA, USA, 2001. ISBN 1-886411-48-4. 396 pp. LCCN QA76.76.O63 S788 2000. US$29.95. URL http:/

/dsl.org/cookbook/. Prepared

with TEXinfo.

Stubbings:2016:OHH

[Stu16] Richard Stubbings, editor. An Oral History of Horsmonden: A series of personal recollections in their own words by villagers born in the early to mid 20th century. Horsmonden Historical Society, Horsmonden, Kent, UK, 2016. ISBN 1-5272-6671-0. 312 pp. LCCN ????

URL http://www.historical-
society.horsmonden.org/. In-

terview team leader Thelma Skinner.

Savi:CET91

[SV91] P. Savi and G. Vecchi. Campi elettromagnetici, Temi d’esame svolti. Edizioni Cooperativa Libraria Universitaria Torino, Corso Duca degli Abruzzi, 24, I 10129 Torino, Italy, 1991. Prepared with L^{A}TEX.

Syzmanski:PFP91

[Syz91] Boleslaw K. Syzmanski. Parallel Functioning Programming Languages and Implementation Techniques. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-52243-8. xiv + 414 pp. LCCN QA76.642 .P37 1991. Prepared with L^{A}TEX.

Tajima:CPP89

[Taj89] Toshiki Tajima. Computational Plasma Physics with Applications to Fusion and Astrophysics. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-16411-6. xxii + 503 pp. LCCN QC718 .T12 1989. US$40.95. Prepared with TEX.

Taylor:CAD91

[Tay91] Dean Taylor. Computer-Aided Design. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-16891-X. xviii + 492 pp. LCCN TA174.T39 1992. Prepared with TEX.

Tesanovic:FTC90

[Tes90] Zlatko Tesanovic, editor. Field Theories in Condensed Matter Physics: A Workshop. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-50391-3. viii + 208 pp. LCCN QC173.4.C65 F54 1990. Prepared with TEX.

Thisted:ESC88

[Thi88] Ronald A. Thisted. Elements of Statistical Computing: Numerical Computation. Chapman and Hall, Ltd., London, UK, 1988. ISBN 0-412-01371-1. 427 pp. LCCN QA276.4.T47 1988. US$45.00. Prepared with TEX.

Thomas:PAG88

[Tho88] Carol G. Thomas. Paths From Ancient Greece. Brill, Leiden, The Netherlands, 1988. ISBN 90-04-08846-6. 206 pp. LCCN CB245 .P336 1988.

Thompson:TTF91

[Tho91] Simon Thompson. Type Theory and Functional Programming. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-41667-0. v + 372 pp. LCCN QA76.62 .T46 1991. Prepared with L^{A}TEX, BibTEX, and makeindex on Apple Macintosh SE/30 using TEXtures, and on a Sun 4. Source files provided to the publisher who modified them to conform to their house style. Page proofs made on laser printer by publisher.

Tick:PLP91

[Tic91] Evan Tick. Parallel Logic Programming. The MIT Press, Cambridge, MA, 1991. ISBN 0-262-20087-2. 486 pp. LCCN QA76.642 .T53 1991. Prepared with L^{A}TEX.

Tondo:MM92

[TNY92] Clovis L. Tondo, Andrew Nathanson, and Eden Yount. Mastering MAKE—A Guide to Building Programs on DOS and UNIX Systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-554619-2. x + 143 pp. LCCN QA76.76.U84 T66 1992. Prepared with L^{A}TEX.

Tondo:MM94

[TNY94] Clovis L. Tondo, Andrew Nathanson, and Eden Yount. Mastering Make—A Guide to Building Programs on DOS, OS/2, and UNIX Systems. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1994. ISBN 0-13-121906-5. xvi + 310 pp. LCCN QA76.76.U84T65 1994. US$22.00. Prepared with L^{A}TEX.

Torchinsky:RV88

[Tor88] Alberto Torchinsky. Real Variables. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-15675-X. xii + 403 pp. LCCN QA331.5.T588 1988. Prepared with TEX.

Troen:AEV90

[TP90a] Ib Troen and Erik Lundtang Petersen. Atlante Europeo del Vento. Risø National Laboratory, Roskilde, Denmark, 1990. ISBN 87-550-1639-1. 264 pp. Typeset with Bitstream fonts on a Canon Series III 300-dpi laser printer, with extensive graphics and data tables; the data files are available on IBM PC diskettes.

Troen:AEE90

[TP90b] Ib Troen and Erik Lundtang Petersen. El Atlas Eólico Europeo. Risø National Laboratory, Roskilde, Denmark, 1990. ISBN 87-550-1638-3. 260 pp. Typeset with Bitstream fonts on a Canon Series III 300-dpi laser printer, with extensive graphics and data tables; the data files are available on IBM PC diskettes.

Troen:EW90

[TP90c] Ib Troen and Erik Lundtang Petersen. Europäischer Windatlas. Risø National Laboratory, Roskilde, Denmark, 1990. ISBN 87-550-1636-7. 240 pp. Typeset with Bitstream fonts on a Canon Series III 300-dpi laser printer, with extensive graphics and data tables; the data files are available on IBM PC diskettes.

Troen:AEE91

[TP91] Ib Troen and Erik Lundtang Petersen. Atlas Éolien Européen. Risø National Laboratory, Roskilde, Denmark, 1991. ISBN 87-550-1637-5. 278 pp. Typeset with Bitstream fonts on a Canon Series III 300-dpi laser printer, with extensive graphics and data tables; the data files are available on IBM PC diskettes.

Turner:RTF90

[Tur90] David A. Turner, editor. Research Topics in Functional Programming. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-17236-4. xv + 374 pp. LCCN QA76.62 .R47 1990. Prepared with L^{A}TEX.

TVguide

[TVG87] TV Guide magazine, 1987. Portions of this magazine (20 million issues weekly) are set with TEX.

Traub:IBC88

[TWW88] J. F. Traub, G. Wasilkowski, and H. Woźniakowski. Information-Based Complexity. Academic Press, Boston, MA, USA, 1988. ISBN 0-12-697545-0. xiii + 523 pp. LCCN QA267 .T731 1988. This book was prepared with AMS-TEX.

Ulichney:DH87

[Uli87]  Robert Ulichney.  Digital

Halftoning. The MIT Press, Cambridge, MA, 1987. ISBN 0-262-21009-6. xiv + 362 pp. LCCN T385 .U451 1987.

UMAP

[UMA] Paul J. Campbell (Beloit College, 700 College St., Beloit, WI 53511-5595, USA), editor. The UMAP Journal. ISSN 0197-3622. COMAP, Inc., Suite 210, 57 Bedford St., Lexington, MA 02173, USA. Prepared with L^{A}TEX. (TeXtures on Apple Macintosh II). Publication with L^{A}TEX began with volume 11 (1990). The Journal is a quarterly (with annual supplement volume) devoted to undergraduate mathematics and its applications, particularly mathematical modeling. It features both articles and instructional modules. A L^{A}TEX style file is available from the editor.

VanDam:FPP89

[Van89] James W. Van Dam, editor.

From Particles to Plasmas: Lectures Honoring Marshall N. Rosenbluth. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-15680-6. xi + 368 pp. LCCN QC717.6 .F76 1987. Prepared with TEX.

vanHerwijnen:1990:PS

[van90] Eric van Herwijnen. Practical SGML. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0635-X, 0-7923-9434-8. xviii + 307 pp. LCCN QA76.73.S44 V36 1990. UK£24.90, US$49.00.

VanLoan:CFF92

[Van92] Charles F. Van Loan. Computational Frameworks for the Fast Fourier Transform. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-285-8. xiii + 273 pp. LCCN QA403.5 .V35 1992. Prepared with L^{A}TEX. Used TEXtures on an Apple Macintosh with private style files.

Varian:IM87

[Var87] Hal R. Varian. Intermediate Microeconomics. W. W. Norton & Co., New York, NY, USA, 1987. ISBN 0-393-95554-0. xviii + 587 pp. LCCN HB172 .V34 1987. Prepared with TEX.

Varian:IM90

[Var90a] Hal R. Varian. Intermediate Microeconomics. W. W. Norton & Co., New York, NY, USA, second edition, 1990. ISBN 0-393-95924-4. xviii + 599 pp. LCCN HB172 .V34 1990. Prepared with TEX.

Varian:M90

[Var90b] Hal R. Varian. Microeconomia. Libreria Editrice Cafoscarina, Venice, Italy, second edition, 1990. ISBN 88-85613-20-9. Prepared with TEX.

Varian:MA91

[Var91] Hal R. Varian. Microeconomic Analysis. W. W. Norton & Co., New York, NY, USA, third edition, 1991. ISBN 0-393-95735-7. xv + 506 pp. LCCN HB172 .V35 1992. Prepared with TEX.

Varian:IM92

[Var92] Hal R. Varian. Intermediate Microeconomics—A Modern Approach. W. W. Norton & Co., New York, NY, USA, third edition, 1992. ISBN 0-393-96320-9. xxiv + 623 pp. LCCN HB172 .V34 1993. Prepared with TEX.

Varian:EFMM93

[Var93] Hal R. Varian, editor. Economic and Financial Modeling with Mathematica. TELOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, first edition, 1993. ISBN 0-387-97882-8 (New York), 3-540-97882-8 (berlin). xxii + 458 pp. LCCN HB143 .E36 1993. Primarily Mathematica Notebooks run through nb2tex conversion software.

vanHerwijnen:PS94

[vH94]  Eric van Herwijnen.  Prac-

tical SGML. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, second edition, 1994. ISBN 0-7923-9434-8. xx + 288 pp. LCCN QA76.73.S44 V36 1994.

Vitter:DAC87

[Vit87] Jeffrey Scott Vitter. Design and Analysis of Coalesced Hashing. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1987. ISBN 0-19-504182-8. xii + 160 pp. LCCN QA76.9.H36 V57 1987. US$29.95. Prepared with TEX.

vanLint:CC92

[vLW92] J. H. van Lint and R. M. Watson. A Course in Combinatorics. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41057-6 (hardcover), 0-521-42260-4 (paperback). ix + 530 pp. LCCN QA164 .L56 1992. Prepared with AMS-TEX.

vanMill:OPT90

[vMR90] Jan van Mill and George M. Reed, editors. Open Problems in Topology. North-Holland Publishing Company, Amsterdam, The Netherlands, 1990. ISBN 0-444-88768-7. xiv + 692 pp. LCCN QA611 .O56 1990. Prepared with L^{A}TEX. K. P. Hart developed a L^{A}TEX style for this book modelled on North-Holland conventions, and developed a font for the pictures in the paper on knot theory.

VAX

[VMS88] VAX VMS version 4.x and 5.x manuals. Digital Equipment Corporation, 1988. The complete VMS manuals sets are produced with TEX, but authors actually prepare input in a form suitable for an earlier in-house system, which is then automatically translated to TEX form.

vonNeumann:2018:MFQd

[vN18] John von Neumann. Mathematical Foundations of Quantum Mechanics. Princeton University Press, Princeton, NJ, USA, 2018. ISBN 0-691-17856-9 (hardcover), 0-691-17857-7 (paperback), 1-4008-8992-8 (e-book). xviii + 304 pp. LCCN QC174.3 .V613 2018.

URL https://www.jstor.org/
stable/j.ctt1wq8zhp. Trans-

lated from the German by Robert T. Beyer. Edited by Nicholas A. Wheeler.

Volid:1992:CBB

[Vol92] Mikhail Volid. Chem byl bolen V. I. Lenin? : meditsinskii i politicheskii analiz (Russian): [What ailed V. I. Lenin? Medical and political analysis]. Effect Publishing, 501 Fifth Avenue, Suite 1612, New York, NY 10017, USA, 1992. ISBN 0-911971-76-9. 76 pp. LCCN DK254.L46 V635 1992.

Wolf:MLT86

[W^{+}86] Kurt Bernardo Wolf et al. Manual de Lenguaje y Tipografía Cientìfica en Castellano. Editorial Trillas, México DF, México, 1986.

Wright:2016:DCG

[WAD16] Christopher Wright, Maria Argyrou, and Charalambos Dendrinos. A Descriptive Catalogue of the Greek Manuscript Collection of Lambeth Palace Library. Lambeth Palace Library and the Hellenic Institute, Royal Holloway, University of London, London, UK, 2016. 544 pp. URL

https://www.royalholloway.
ac.uk/Hellenic-Institute/
Research/LPL/Greek-MSS/Catalogue.
pdf.

Wakerly:MAP89

[Wak89] John F. Wakerly. Microcomputer architecture and programming: the 68000 family. John Wiley and Sons, Inc., New York, NY, USA, 1989. ISBN 0-471-85319-4. xxii + 762 pp. LCCN QA76.8.M6895 W35 1989. Prepared with L^{A}TEX. Includes bibliographies and index.

Wattenberg:PMC90

[Wat90] Frank Wattenberg. Personal Mathematics and Computing. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-23157-3. xiv + 556 pp. LCCN QA39.2.W378 1990.

Winston:AI79

[WB79] Patrick Henry Winston and R. H. Brown, editors. Artificial Intelligence: An MIT Perspective (2 volumes). The MIT Press, Cambridge, MA, 1979. ISBN 0-262-23096-8. LCCN Q335 .A7865 1979. Prepared with TEX.

Wei:TSA90

[Wei90a] William Wei. Time Series Analysis: Univariate and Multivariate Methods. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-15911-2. xv + 478 pp. LCCN QA280 .W45 1990. Prepared with TEX.

Weir:GDA90

[Wei90b] Bruce S. Weir. Genetic Data Analysis. Sinauer Associates, Inc., Sunderland, MA 01375, USA, 1990. ISBN 0-87893-871-0 (hardcover), 0-87893-872-9 (paperback). xii + 377 pp. LCCN QH438.4.S73 W45 1990. US$48.00 (hardcover), US$27.00 (paperback).

Weiss:IS91

[Wei91] Neil Weiss. Introduction to Statistics. Addison-Wesley, Reading, MA, USA, third edition, 1991. ISBN 0-201-17833-8. xviii + 834 pp. LCCN QA276.12 .W45 1991. Prepared with TEX.

Weisstein:1999:CCE

[Wei99] Eric W. Weisstein. The CRC Concise Encyclopedia of Mathematics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-9640-9. 1969 pp. LCCN QA5.W45 1999. US$79.95.

Wellcome:MMII89

[Wel89a] Wellcome Institute for the History of Medicine. A Mirror for Medicine: II. Some resources of the Wellcome Institute. An Exhibition Monday 12 June— Friday 29 September 1989. Wellcome Institute for the History of Medicine, London, UK, 1989. 42 pp. LCCN WZ 23 W4475w 1989. Prepared with L^{A}TEX. PCTEX and Computer Modern fonts, author. It was based on the style file that he wrote for Traub et al., Information-Based Complexity, with various vital parameters being lifted from Oxford’s inhouse L^{A}TEX style file.

Wellcome:MMI87

[Wel89b] Wellcome Institute for the History of Medicine. A Mirror for Medicine: Some resources of the Wellcome Institute. An Exhibition Monday 19 October— Friday 18 December 1987. Wellcome Institute for the History of Medicine, London, UK, 1989. Prepared with L^{A}TEX. PCTEX and Computer Modern fonts, with proofed on HP LaserJet II at 300dpi, then converted to PostScript using James Clark’s dvitops, with final output on bromide at 1270dpi on the ULCC Linotronic 300.

Weld:TCA90

[Wel90] Daniel Weld. Theories of Comparative Analysis. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-23152-2. xii + 184 pp. LCCN Q335 .W42 1990.

Werschulz:CCD91

[Wer91] Arthur G. Werschulz. The Computational Complexity of Differential and Integral Equations: An Information-Based Approach. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, October 1991. ISBN 0-19-853589-9. ix + 331 pp. LCCN QA372 .W47 1991. US$52.50. Prepared with AMS-TEX. The style file was written by the final output on an HP LaserJet II at 300dpi.

Waser:IAD82

[WF82] Shlomo Waser and Michael J. Flynn. Introduction to Arithmetic for Digital Systems Designers. Holt, Rinehart, and Winston, New York, NY, USA, 1982. ISBN 0-03-060571-7. xvii + 308 pp. LCCN TK7895 A65 W37 1982. Prepared with TEX. Master copy output on Alphatype CRS high-resolution phototypesetter. This book went to press while the IEEE 754 Floating-Point Standard was still in development; consequently, some of the material on that system was invalidated by the final Standard (1985).

Weiss:PWR93

[WG93] Zbigniew Weiss and Tadeusz Gruźlewski. Programowanie współbieżne i rozproszone w przykładach i zadaniach. Wydawnictwa Naukowo-Techniczne, Warszawa, Polska, 1. edition, 1993. ISBN 83-204-1637-X. składał: Zbigniew Weiss, druk: Drukarnia Wydawnictw Naukowych s.a. Łódź, ul. Żwirki 2.

Winston:L88

[WH88] Patrick Henry Winston and Berthold Klaus Paul Horn. Lisp. Addison-Wesley, Reading, MA, USA, third edition, 1988. ISBN 0-201-08319-1. xxi + 611 pp. LCCN QA76.73.L23 W56 1989. Prepared with TEX.

Winston:L81

[WH89] Patrick Henry Winston and Berthold Klaus Paul Horn. Lisp. Addison-Wesley, Reading, MA, USA, 1981, 1984, 1989. ISBN 0-201-08319-1. xxi + 611 pp. LCCN QA76.73.L23 W56 1989. First edition may be one of the first books published using TEX. Macros used there, developed by Daniel Brotsky, later formed the basis of what is now YTEX.

Ward:CS90

[WH90] Stephen A. Ward and Robert H. Halstead Jr. Computation Structures. The MIT Electrical Engineering and Computer Science series. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-23139-5 (MIT), 0-07-068147-3 (McGraw-Hill). xx + 789 pp. LCCN TK7888.4 .W37 1990. Copublished by McGraw-Hill, New York.

Bolch:LR89

[wHR89] Gunter Bolch with Helmut Riedel. Leistungsbewertung von Rechensystemen. B. G. Teubner, Stuttgart, Germany, 1989. ISBN 3-519-02279-6. Prepared with L^{A}TEX.

Wiederhold:FOD87

[Wie87] Gio Wiederhold. File Organization for Database Design. McGraw-Hill Computer Science Series, Editor: Richard W. Hamming and Edward A. Feigenbaum. McGraw-Hill, New York, NY, USA, second edition, 1987. ISBN 0-07-070133-4. 619 pp. LCCN QA76.9.F5 W53 1987. US$38.95.

Wiggins:IAN90

[Wig90] Stephen Wiggins. Introduction to Applied Nonlinear Dynamical Systems. R. R. Donnelley and Sons, Harrisburg, VA, USA, 1990. ISBN 3-540-97003-7. xiv + 672 pp. LCCN QA614.8 .W54 1990. Prepared with L^{A}TEX.

Wilf:AC86

[Wil86] Herbert S. Wilf. Algorithms and Complexity. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-021973-8 (hardcover), 0-13-022054-X (paperback). vi + 231 pp. LCCN QA63 .W55 1986.

Winder:CTM89

[Win89a] Marianne Winder. Catalogue of Tibetan Manuscripts and Xylographs, and Catalogue of Thankas, Banners and other Paintings and Drawings in the Library of the Wellcome Institute for the History of Medicine. Wellcome Institute for the History of Medicine, London, UK, 1989. ISBN 0-85484-085-0. Prepared with L^{A}TEX. PCTEX and Computer Modern fonts, with final output on an HP LaserJet II at 300dpi.

Winston:AI89

[Win89b] Patrick Henry Winston. Artificial Intelligence. Addison-Wes-ley, Reading, MA, USA, second edition, 1989. ISBN 0-201-09696-X. Prepared with TEX.

Wujastyk:MLD86

[wJBH86] Dominik Wujastyk with Judith Barker and Katy Hooper. A Manual of Library Desk Practice as Practised at the Wellcome Institute for the History of Medicine . . .. Wellcome Institute for the History of Medicine, London, UK, 1986. Prepared with L^{A}TEX. PCTEX and Computer Modern fonts, with final output on a Canon LBP8 A1 at 300dpi. The printer was managed by software from Personal TEX driving a Flagstaff Engineering JLaser card plugged into the PC bus, and sending bitmaps directly to the laser printer, bypassing the controller.

Woodcock:SEM89

[WL89] Jim Woodcock and Martin Loomes. Software Engineering Mathematics. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-50424-3. xii + 271 pp. LCCN QA76.758 .W66 1989. Prepared with L^{A}TEX.

Wolf:NP83

[Wol83] Kurt Bernardo Wolf, editor. Nonlinear Phenomena, volume 189 of Lecture Notes in Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-387-12730-5, 3-540-12730-5. x + 451 pp. LCCN QC20.7.N6 S36 1982. DM58.00. First book in TEX published by Springer-Verlag.

Wolfram:MSD88

[Wol88] Stephen Wolfram. Mathematica: A System for Doing Mathematics by Computer. Addison-Wes-ley, Reading, MA, USA, 1988. ISBN 0-201-19334-5 (hardcover), 0-201-19330-2 (paperback). xviii + 749 pp. LCCN QA76.95 .W65 1988. US$29.95. This book was prepared with TEX, L^{A}TEX, and PostScript.

Wolf:LMO89

[Wol89] Kurt Bernardo Wolf, editor. Lie Methods in Optics, vol. II, volume 352 of Lecture Notes in Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-52123-2, 0-387-52123-2. xii + 197 pp. LCCN QC355.2 .L54 1988. US$29.70.

Wolfram:MSD90

[Wol90a] Stephen Wolfram. Mathematica:

A System for Doing Mathematics by Computer. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-19334-5. xviii + 749 pp. LCCN QA76.95 .W65 1988. Prepared with TEX.

Mathematica

[Wol90b] Stephen Wolfram. Mathematica journal. Addison-Wesley, 1990. ISSN 1047-5974.

Winston:AIB84

[WP84] Patrick H. Winston and Karen Prendergast, editors. The A.I. Business: The Commercial Uses of Artificial Intelligence. The MIT Press, Cambridge, MA, 1984. ISBN 0-262-23117-4. 324 pp. LCCN Q334 .A45 1984. Prepared with TEX.

Winston:AIM90-1

[WS90a] Patrick Winston and Sarah Alexandra Shellard, editors. Artificial Intelligence at MIT, volume 1. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-23150-6 (volume two), 0-262-23154-9 (two-volume set). LCCN Q335.A78713 1990.

Winston:AIM90-2

[WS90b] Patrick Winston and Sarah Alexandra Shellard, editors. Artificial Intelligence at MIT, volume 2. The MIT Press, Cambridge, MA, 1990. ISBN 0-262-23151-4 (volume two), 0-262-23154-9 (two-volume set). LCCN Q335.A78713 1990.

Watt:ALM87

[WWF87] David A. Watt, Brian A. Wichmann, and William Findlay. Ada Language and Methodology. Prentice-Hall international series in computer science. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-004078-9 (paperback), 0-13-004086-X (paperback with disk). xvii + 518 pp. LCCN QA76.73.A35 W38 1987. US$28.67. Prepared with TEX (unknown variety).

Zurek:CEP90

[Zur90] Wojciech Zurek, editor. Complexity, Entropy and the Physics of Information. Addison-Wes-ley, Reading, MA, USA, 1990. ISBN 0-201-51509-1 (hardcover), 0-201-51506-7 (paperback). xiii + 530 pp. LCCN QC39.S48 1991. Prepared with TEX.

Zwillinger:HDE89

[Zwi89] Daniel Zwillinger. Handbook of Differential Equations. Academic Press, Boston, MA, USA, 1989. ISBN 0-12-784390-6. xx + 673 pp. LCCN QA371 .Z88 1989.

Zwillinger:1992:HI

[Zwi92] Daniel Zwillinger. Handbook of Integration. A. K. Peters, Ltd., Wellesley, MA, USA, 1992. ISBN 0-86720-293-9, 1-4398-6584-1 (ebook). xv + 367 pp. LCCN QA299.3 .Z85 1992. Prepared with TEX.